### v2

# Parameter-Efficient Subspace Optimization for LLM Fine-Tuning

Abstract
This paper develops a new perspective on parameter-efficient fine-tuning (PEFT) for LLMs, inspired by classical subspace minimization. We introduce a unifying framework, Parameter-Efficient Subspace Optimization (PESO), which recovers existing methods such as LoRA and connects them to the principled algorithmic and theoretical foundations of subspace optimization. This connection highlights a natural “exploration–exploitation” view of subspace methods, guiding the design of new algorithms that achieve strong convergence performance while still preserving memory efficiency. We instantiate the framework into a practical algorithm, PESO-LoRA, based on a LoRA-type parameterization. Importantly, we provide convergence guarantees stated in the full-parameter space for the induced update, addressing a key limitation of LoRA-style analyses that only track low-dimensional factors. Empirically, PESO-LoRA improves over strong PEFT baselines on standard fine-tuning benchmarks.
1 Introduction
Training deep neural networks is the cornerstone of modern AI, powering the success of large-scale foundation models such as Large Language Models (LLMs) (Brown et al., 2020). At the core, it reduces to solving a high-dimensional optimization problem over weight matrices:
$\Delta W^{*}:=\arg\min_{\Delta W}\,\,\ell\big(W_{0}+\Delta W\big),$ (1)
where $\ell(\cdot)$ is the loss function, $W_{0}$ is the initialization, and $\Delta W$ the increment. In practice, (1) is typically solved by first-order methods such as Adam (Kingma and Ba, 2014) and AdamW (Loshchilov and Hutter, 2017), which are the workhorses of large-scale training. However, these methods require storing additional optimizer states (e.g., momentum and velocity), and for LLMs this overhead places enormous pressure on memory resources, making parameter-efficient strategies appealing.
In the realm of fine-tuning, we often have limited labeled data for a downstream task but still wish to adapt the pretrained weights effectively and efficiently. Therefore, updating the entire parameter set is memory-intensive. This motivates the study of Parameter-Efficient Fine-Tuning (PEFT) methods (Han et al., 2024; Houlsby et al., 2019; Hu et al., 2022), where optimization is restricted to a smaller set of parameters initialized from pretrained weights. In other words, $W_{0}$ denotes weights obtained from a large-scale pretraining phase, and $\Delta W$ is not updated freely but instead follows an efficient parameterization that constrains the search space.
Figure 1: Comparison of full-parameter tuning, LoRA, and our method (PESO-LoRA). Left: optimization trajectories. Top Right: MetaMathQA. Bottom Right: synthetic example $\min_{W}\|W-M\|_{F}^{2}$ with $M=10\cdot\operatorname{diag}(1,\dots,1,0,\dots,0)$ (r+1r{+}1 ones); see Appendix A. PESO-LoRA bridges the loss gap of LoRA while preserving memory and computation efficiency.
A popular PEFT method is low-rank adaptation (LoRA, Hu et al. (2022)), where matrices in $\Delta W$ are expressed as the product of two low-rank factors. LoRA has shown strong empirical success, reducing memory costs while achieving competitive downstream performance. However, it suffers from two key limitations: 1) performance often lags behind full-parameter fine-tuning (Figure 1, top right: MetaMathQA); 2) theoretical guarantees are limited, with convergence typically shown only for the low-rank factors (Figure 1, bottom right: a synthetic example illustrating LoRA’s potentially unbounded loss gap). To address these issues, many LoRA variants (Hayou et al., 2024; Wang et al., 2024a, b; Zhang et al., 2023, 2025b) have been proposed, yet they largely inherit the same shortcomings and leave the following fundamental question open:
Can we design fine-tuning methods that maintain the practical performance of LoRA while still enjoying the convergence and optimality of full-parameter fine-tuning?
To address this question, we reveal an inherent connection between PEFT and the classical idea of subspace minimization, a long-standing nonlinear optimization strategy dating back to Conn et al. (1994); Cragg and Levy (1969). The central philosophy is to decompose a large-scale problem like (1) into iterative, simpler subproblems constrained to carefully chosen subspaces. This view resonates naturally with modern PEFT methods, which restrict updates to structured low-rank forms for better efficiency. Interestingly, subspace minimization historically received less attention in the optimization society, since full-parameter information were often affordable in traditional applications. However, it is especially well suited to LLM training, where massive dimensionality calls for memory-efficient methods.
Formally, we build on the notion of intrinsic dimensionality in LLM training (Aghajanyan et al., 2020; Li et al., 2018), recognized in (Hu et al., 2022) as the origin of LoRA: there exists a dimension-lifting map $\mathcal{M}:\mathbb{R}^{d}\to\mathbb{R}^{m\times n}$, with $d\ll m\times n$, such that the optimal solution $\Delta W^{*}$ of (1) satisfies
$\Delta W^{*}\approx\mathcal{M}(\xi^{*}),\,\xi^{*}:=\arg\min_{\xi\in\mathbb{R}^{d}}\ell(W_{0}+\mathcal{M}(\xi)).$ (2)
Here, $d$ stands for the number of trainable parameters, and this characterization implies that it suffices to optimize within the reduced space defined by the image of $\mathcal{M}$ to approximate $\Delta W^{*}$. For clarity, we focus on a single weight matrix $\Delta W\in\mathbb{R}^{m\times n}$ (multi-layer extensions are straightforward) and represent $\xi$ as a $d$-dimensional vector. This is without loss of generality, since tensor parameters can always be flattened via vectorization into an isomorphic Euclidean space. For example, LoRA adopts the simple form $\mathcal{M}(A,B)=AB$ with $A\in\mathbb{R}^{m\times r}$, $B\in\mathbb{R}^{r\times n}$, and $d=(m+n)r$. However, it remains unclear whether such a simple $\mathcal{M}$ is sufficient to capture the complexity of LLM training dynamics.
Our framework approximates $\mathcal{M}$ adaptively through a sequential subspace approximation, providing a more effective capture of (2). We construct a sequence of maps $\{\mathcal{M}_{k}\}$, each with a simple representation and $\mathbb{R}^{d}\to\mathbb{R}^{m\times n}$,
$\begin{array}[]{l}\Delta W^{*}\approx\sum_{k}\mathcal{M}_{k}(\xi_{k}^{*}),\\ \xi_{k}^{*}:=\arg\min_{\xi}\ell\!\left(W_{0}+\sum_{i=1}^{k-1}\mathcal{M}_{i}(\xi_{i}^{*})+\mathcal{M}_{k}(\xi)\right).\end{array}$ (3)
Each image of $\mathcal{M}_{k}$ approximates a subspace and $\xi_{k}^{*}$ is its low-dimensional coordinate. In essence, the complexity of $\mathcal{M}$ is captured by a sequence of piecewise-linear subspaces. This philosophy parallels classical approximation schemes in numerical analysis such as finite element methods (Bathe, 2006).
Guided by this perspective, we develop a principled framework for PEFT grounded in subspace minimization, named Parameter-Efficient Subspace Optimization (PESO). A key insight is to view the problem (3) through an exploration–exploitation lens: exploration designs new subspaces that capture full gradient information, while exploitation optimizes efficiently within the current subspace. This resolves LoRA’s two central limitations: lack of full-parameter convergence guarantees and inefficiency from rigid low-rank parameterization; see Figure 1.
Contributions. Our contributions can be summarized at three levels. Although our focus is on PEFT, many of the ideas developed here naturally extend to pre-training.
I. Perspective Level.
We introduce a novel framework PESO for memory-efficient training inspired by classical subspace minimization (Conn et al., 1994), unifying existing PEFT approaches such as LoRA variants (Hu et al., 2022; Wang et al., 2024a, b; Zhang et al., 2023, 2025b) and GaLore (Zhao et al., 2024). This framework allows us to explore the rich algorithmic techniques in subspace methods, providing systematic guidance to improve memory-efficient methods. In particular, we highlight two complementary directions: exploration of new subspaces through information from the full gradient, and exploitation of the current subspace via streaming SVD representations.
II. Theoretical Level.
Our exploration mechanism, full gradient restart, enables the framework to effectively guide training dynamics. The resulting algorithm is tailored to LLM fine-tuning and combines the empirical efficiency of LoRA-style designs with convergence guarantees stated in the full-parameter space for the induced update. This aspect is largely absent from existing LoRA-style analyses, which typically track only the low-dimensional factors.
III. Empirical Level.
Guided by our PESO framework, we show that two practical instantiations of our framework—PESO-LoRA-R and PESO-LoRA-T—achieve improved performance while preserving the memory efficiency of state-of-the-art PEFT methods across benchmarks such as GLUE, mathematical reasoning, code generation, and general instruction tuning.
Related Work. LoRA (Hu et al., 2022) is perhaps the most widely known PEFT method, and numerous variants have been proposed for better performance. For instance, LoRA+ (Hayou et al., 2024) introduces imbalanced learning rates; PiSSA (Meng et al., 2024) proposes an initialization from SVD of $W_{0}$; and AdaLoRA (Zhang et al., 2023) maintains an adaptive SVD-based low-rank representation. Other extensions focus on gradient scaling (Tastan et al., 2025; Zhang and Pilanci, 2024). More recent work leverages information from the full gradient: LoRA-GA (Wang et al., 2024a) and LoRA-Pro (Wang et al., 2024b) propose memory-efficient gradient approximations, and LoRA-One (Zhang et al., 2025b) employs the SVD of the full gradient for initialization.
One closely related line of work studies memory-efficient training via projection or subspace updates in the pre-training regime. GaLore (Zhao et al., 2024) constructs the projection subspace using a top-$r$ SVD of the full gradient, while Fira (Chen et al., 2024) augments this with a gradient-correction step to reduce projection bias. SARA (Zhang et al., 2025a) employs importance sampling to choose subspaces based on the SVD spectrum of the full gradient. APOLLO (Zhu et al., 2024) instead uses randomized projections to define the subspace, and SubTrack++ (Rajabi et al., 2025) explores Grassmannian subspaces for improved tracking. While these pre-training methods can be adapted to PEFT, they are not designed to exploit the strong intrinsic dimensionality (2) in PEFT tasks and can underperform LoRA-style parameterizations; see Table 5. In contrast, PESO builds on classical subspace minimization (Conn et al., 1994), embedding LoRA-type structures and viewing fine-tuning as sequential subproblems that iteratively capture intrinsic dimensionality (3).
Convergence guarantees for PEFT algorithms remain scarce, and existing results typically address only the low-dimensional parameters (Jiang et al., 2024). A closely related line of work studies randomized subspace descent (Kozak et al., 2019), including RSO (Chen et al., 2025) and GoLore (He et al., 2024), which construct subspaces via randomized projections and provide convergence guarantees in the full-parameter space. Their exploration strategies differ from ours and they typically do not encode PEFT-specific LoRA structures. Other approaches require additional structural conditions, e.g., Liang et al. (2024) assumes the projection subspace has full column rank (unrealistic when $r<m$) and LDAdam (Robert et al., 2024) assumes a strict contraction. Closer to LoRA, Jang et al. (2024) analyze LoRA only in the NTK regime, and RAC-LoRA (Malinovsky et al., 2024) proposes a randomized asymmetric-chaining variant, a different exploration mechanism from ours.
Subspace minimization is a classical theme in nonlinear optimization (Conn et al., 1994; Cragg and Levy, 1969; Yuan, 2014). It was historically overshadowed by full-parameter algorithms such as L-BFGS (Liu and Nocedal, 1989) and conjugate gradient methods (Nocedal and Wright, 2006, Ch. 5), since many traditional applications could afford storing full gradients and quasi-Newton pairs. More recently, however, subspace-based strategies have re-emerged in large-scale derivative-free optimization (Cartis and Roberts, 2023; Dzahini and Wild, 2024; Menickelly, 2024; Nozawa et al., 2025; Zhang, 2025), where gradients are unavailable and low-dimensional surrogates are crucial.
2 PESO: A Framework From Subspace Minimization
In this section, we provide a novel perspective of PEFT methods with insights from subspace minimization. We summarize an algorithmic framework Parameter-Efficient Subspace Optimization (PESO) in Algorithm 1, and discuss how it unifies many benchmarks such as LoRA and GaLore.
To build an iterative scheme, a central question is how to represent the weight $W$ at each iteration using low-dimensional representation $\xi$. In (3), the optimization is expressed through evolving subspaces. At iteration $k$, we define the anchored state $\widetilde{W}_{k}:=W_{0}+\sum_{i=1}^{k-1}\mathcal{M}_{i}(\xi_{i}^{*})$ to encode historical progress, and represent
$W_{k}=\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k}).$ (4)
Following the design of subspace minimization, PESO considers each $\mathcal{M}_{k}$ to admit a simple image in the form of a subspace: $\mathcal{S}_{k}:=\{\mathcal{M}_{k}(\xi):\xi\in\mathbb{R}^{d}\}$.
Under representation (4), the evolution of $W_{k}$ can be viewed as three complementary operations: 1) exploration: updating $\mathcal{M}_{k}$ to select a new subspace $\mathcal{S}_{k}$, (line 6 of Algorithm 1) 2) exploitation: optimizing $\xi_{k}$ within the current subspace (line 8-9 of Algorithm 1), and 3) updating $\widetilde{W}_{k}$ to absorb past progress into the anchored weights. These operations mirror the classical paradigm of subspace minimization (Conn et al., 1994), where a large-scale problem is solved by iteratively: (i) constructing a subspace based on local information such as gradients, (ii) solving a reduced subproblem within that subspace, (iii) updating the iterate to incorporate the subspace solution.
In our design, exploration and exploitation directly parallel subspace selection and subproblem optimization, while the anchored state $\widetilde{W}_{k}$ retains progress from earlier subspaces. In LoRA, $\widetilde{W}$ is fixed at $W_{0}$, confining progress to the active subspace. In contrast, updating $\widetilde{W}$ absorbs accumulated contributions back into the parameter space, giving rise to two distinct exploration strategies: warm-start and restart, which we detail below in Section 2.1.
Leveraging this connection to subspace minimization, we present our generic framework PESO in Algorithm 1. It is important to note that, by selecting corresponding parameterization of $\mathcal{M}_{k}$, UpdateM, and Opt, we are able to recover a variety of existing benchmarking methods in parameter-efficient training; see representatives in Table 1. We also remark that Algorithm 1 is equivalent to the classical two-loop subspace minimization scheme (Conn et al., 1994), which we defer to Appendix B in Algorithm 4.
Algorithm 1 PESO: Generic Framework of Parameter-Efficient Subspace Optimization
1:Initialization $W_{0}\in\mathbb{R}^{m\times n}$, $\xi_{0}\in\mathbb{R}^{d}$, and $\mathcal{M}_{0}$; an algorithmic subroutine UpdateM, an optimizer Opt, frequency $K$.
2:Set $k\leftarrow 1$ and $\widetilde{W}_{0}\leftarrow W_{0}$.
3:while stopping criteria not satisfied do
4:  $(\mathcal{M}_{k},\widetilde{W}_{k})\leftarrow(\mathcal{M}_{k-1},\widetilde{W}_{k-1})$.
5:  if $k-1\bmod K=0$ then $\triangleright$ Explore new $\mathcal{S}_{k}$
6:   $(\mathcal{M}_{k},\widetilde{W}_{k})$ $\leftarrow{\texttt{UpdateM}}(\mathcal{M}_{k-1},\widetilde{W}_{k-1})$.
7:  end if
8:  $\Delta\xi_{k}\leftarrow{\texttt{Opt}}(\xi_{k-1},\mathcal{M}_{k})$ $\triangleright$ Exploit current $\mathcal{S}_{k}$
9:  $\xi_{k}\leftarrow\xi_{k-1}+\Delta\xi_{k}$.
10:  $k\leftarrow k+1$.
11:end while
2.1 Subspace Exploration-Exploitation in PESO
Now let us discuss two main components of our framework, subspace exploration and exploitation.
Subspace Exploration.
Exploring new subspaces is essential for navigating the full-parameter space under memory restriction. Algorithm 1 carries out exploration by UpdateM, which updates both $\mathcal{M}_{k}$ and $\widetilde{W}_{k}$. Such updates are often performed lazily every $K$ iterations, as in prior works (Liang et al., 2024; Zhang et al., 2023; Zhao et al., 2024; Zhu et al., 2024).
Depending on how much $\mathcal{M}_{k}$ is changed, two philosophies arise for how exploration interacts with the low-dimensional $\xi_{k}$: warm-start and restart. These are simply two modes of UpdateM:
  • Warm-start. Preserve $\xi_{k}$ and keep $\widetilde{W}_{k}$ fixed:
    $W_{k+1}=\widetilde{W}_{k}+\mathcal{M}_{k+1}(\xi_{k}+\Delta\xi_{k}).$ (5)
  • Restart. Absorb the previous contribution into the baseline, $\widetilde{W}_{k+1}\leftarrow\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k})$, and start the new subspace from $\xi_{\text{new}}$ (often 0):
    $W_{k+1}=\widetilde{W}_{k+1}+\mathcal{M}_{k+1}(\xi_{\text{new}}+\Delta\xi_{k}).$ (6)
Intuitively, warm-start provides smoother transitions when consecutive subspaces remain similar, while restart marks a new phase, useful when the optimization geometry changes sharply. In practice, these two modes naturally lead to two main approaches for designing UpdateM. A warm-start typically updates the parameterization of $\mathcal{M}_{k}$ smoothly along an optimization trajectory—for example, by applying an Adam step on subspace parameters as in LoRA variants—yielding a gradually evolving subspace. Restart, on the other hand, often reassigns $\mathcal{M}_{k}$ directly using local information such as gradients. This strategy is common in classical optimization; for example, in line search (a one-dimensional subspace method) each iteration resets the step size initialization when a new direction is chosen (Nocedal and Wright, 2006, Ch. 3). It is also used in LLM training, as in GaLore (Zhao et al., 2024), which periodically resets the subspace via the SVD of the full gradient. Concrete examples of both approaches are summarized in Table 1, and Section 3.1 introduces a new restart scheme leveraging full gradients.
Subspace Exploitation.
Between two updates of UpdateM, our framework performs $K$ iterations of Opt within the current subspace $\mathcal{S}_{k}$. This amounts to solving the subproblem
$\min_{\xi\in\mathbb{R}^{d}}\,\,\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi))$ (7)
approximately for $K$ steps. In practice, Opt is often chosen as Adam.
The philosophy relies on a common belief in classical optimization: during an optimization procedure, once an effective subspace is identified, repeatedly exploiting it for multiple iterations improves efficiency. This principle underlies many classical optimization methods, such as trust-region methods (Nocedal and Wright, 2006, Ch. 4) and L-BFGS-B (Byrd et al., 1995).
2.2 Connection to Existing Benchmarks
While (4) may strike to be abstract, many existing benchmarks for LLM training can naturally fit in it by considering specific subspaces. Here we summarize several notable methods in Table 1.
Table 1: Examples of memory-efficient training methods as instances of PESO.
[…]
Projected subspace. A simple way to define a memory-efficient subspace is through low-rank projection, where $\mathcal{M}_{k}:R\in\mathbb{R}^{r\times n}\mapsto P_{k}R$ is parameterized by a left-projection matrix $P_{k}\in\mathbb{R}^{m\times r}$. This formulation can be extended to right-sided or two-sided projections. By applying the chain rule to $\nabla_{R}\ell(\widetilde{W}_{k}+P_{k}R)$, one obtains the projected subspace schemes analyzed in (He et al., 2024; Kozak et al., 2019; Liang et al., 2024; Zhao et al., 2024); see Appendix C for details. Within PESO, GaLore (Zhao et al., 2024), APOLLO (Zhu et al., 2024), and stochastic subspace descent (Kozak et al., 2019) correspond to a restart strategy by reassigning $P_{k}$, while online subspace descent (Liang et al., 2024) adopts a warm-start update of $P_{k}$ via online PCA.
Low-rank subspace. The LoRA family defines the subspace $\{A_{k}B+AB_{k}:A\!\in\!\mathbb{R}^{m\times r},\,B\!\in\!\mathbb{R}^{r\times n}\}$, where the adapters $(A,B)$ jointly serve as both $\xi$ and the parameterization of $\mathcal{M}_{k}$. Consequently, a single Adam update of $(A,B)$ simultaneously updates the subspace and its coordinates, effectively realizing a warm-start scheme with $K=1$. Many LoRA variants can be viewed as modifications of this generic template: LoRA-Pro (Wang et al., 2024b) applies a different preconditioner, and PiSSA (Meng et al., 2024), LoRA-GA (Wang et al., 2024a), and LoRA-One (Zhang et al., 2025b) adjust initialization strategies, while other works modify learning rates or scaling rules. Our framework unifies these designs by interpreting them as specific choices of Opt or initialization within the same subspace structure.
SVD subspace. A principled way to extract low-dimensional structure from matrices (such as $\Delta W$) is through Singular Value Decomposition (SVD), leading to the representation $\mathcal{M}_{k}:\lambda\in\mathbb{R}^{r}\mapsto U\mathrm{diag}(\lambda)V$. Here, $(U,V)$ define the subspace (exploration), while $\lambda$ is the low-dimensional coordinates (exploitation). This separation fits directly into Algorithm 1, enabling flexible optimization strategies for $(U,V)$ and $\lambda$. AdaLoRA (Zhang et al., 2023) exemplifies this parameterization, and our framework clarifies the roles of $(P_{k},\Lambda_{k},Q_{k})$ in their notation. We build on this principle in Section 3, where we propose a practical SVD-based variant PESO-LoRA-T.
Together, these three subspace categories illustrate how PESO unifies existing PEFT methods under a single framework. This perspective guides the design of practical algorithms with new exploration and exploitation techniques for PEFT, and it sets up the convergence analysis in the sections that follow.
3 PESO-LoRA: Practical Algorithms from the Framework
The previous section establishes a conceptual link between PEFT methods and classical subspace minimization, providing a unifying interpretation. Building on this view, we now develop a concrete algorithm PESO-LoRA, which combines PESO’s optimization principles with LoRA’s empirically effective PEFT parameterization. We present two variants: PESO-LoRA-R leverages a full gradient restart strategy to improve exploration of subspaces, and PESO-LoRA-T is a SVD-based method that enhances exploitation through more effective optimization within each subspace.
3.1 Full Gradient Restart
We now introduce an important variant of the UpdateM subroutine in the restart category (see (6)) that enables convergence to stationarity in the full-parameter space. We design UpdateM so that each new subspace $\mathcal{S}_{k}$ induced by $\mathcal{M}_{k}$ remains well aligned with the current full gradient $G_{k}$. We call this scheme full gradient restart:
Full Gradient Restart. Given learning rates $\{\eta_{k}\}$, whenever $k-1\bmod K=0$:
  1. 1)
    Absorb history: $\widetilde{W}_{k}\leftarrow\widetilde{W}_{k-1}+\mathcal{M}_{k-1}(\xi_{k-1})$.
  2. 2)
    Compute the (stochastic) full gradient $G_{k}=\nabla_{W}\ell(\widetilde{W}_{k})$.
  3. 3)
    Choose a low rank subspace $\mathcal{S}_{k}^{\text{FG}}$ depending on $G_{k}$.
  4. 4)
    Restart with $\xi_{k}\leftarrow\xi_{k}^{\text{new}}$ such that $\mathcal{M}_{k}(\xi_{k}^{\text{new}})=-\eta_{k}P_{\mathcal{S}_{k}^{\text{FG}}}(G_{k})$.
Here, $P_{\mathcal{S}_{k}^{\text{FG}}}(G_{k})$ denotes the projection of $G_{k}$ onto $\mathcal{S}_{k}^{\text{FG}}$. This procedure effectively redefines $\mathcal{M}_{k}$ so that the new adapter is initialized by a projected gradient step:
$W_{k}\leftarrow W_{k-1}-\eta_{k}P_{\mathcal{S}_{k}^{\text{FG}}}(G_{k}).$ (8)
Thus, each restart ensures that $\mathcal{S}_{k}$ captures information from full gradients, with initial progress comparable to a standard SGD step. In the literature on subspace methods, incorporating the full gradient into $\{\mathcal{S}_{k}\}$ is critical for convergence guarantees (Conn et al., 1994; Zhang, 2025). In particular, one can show that $\|\nabla_{W}\ell\|\to 0$ provided that $G_{k}:=\nabla_{W}\ell(W_{k})$ lies in $\mathcal{S}_{k}$. Building on this, we demonstrate in Section 5 that full gradient restart ensures convergence to a stationary point of the original problem (1) by interleaving projected steepest-descent steps with subspace updates.
Algorithm 2 PESO-LoRA-R: PESO with LoRA and Subspace ExploRation
1:Pre-trained parameters $W_{0}\in\mathbb{R}^{m\times n}$, frequency $K$, scale parameter $\gamma$.
2:Set $k\leftarrow 1$, $\widetilde{W}_{0}\leftarrow W_{0}$, $A_{0}\leftarrow 0$ and $B_{0}\leftarrow 0$.
3:while stopping criteria not satisfied do
4:  if $k-1\mod{K}=0$ then
5:   $\widetilde{W}_{k}\leftarrow\widetilde{W}_{k-1}+A_{k-1}B_{k-1}$.
6:   Compute stochastic full gradient $G_{k}$.
7:   $(U_{k},\Lambda_{k},V_{k})\leftarrow{\tt SVD}(-G_{k})$.
8:   $A_{k-1}\leftarrow\frac{1}{\sqrt{\gamma}}U_{k}\sqrt{\Lambda_{k}}$, $B_{k-1}\leftarrow\frac{1}{\sqrt{\gamma}}\sqrt{\Lambda_{k}}V_{k}$.
9:  end if
10:  $(A_{k},B_{k})\leftarrow\texttt{AdamW}(A_{k-1},B_{k-1})$.
11:  $k\leftarrow k+1$.
12:end while
13:return $\widetilde{W}_{k}+A_{k}B_{k}$.
A practical construction of $\mathcal{S}_{k}^{\text{FG}}$ is to compute a rank-$r$ SVD of $G_{k}$ and define the subspace as the span of its top singular directions. This ensures that $\mathcal{S}_{k}^{\text{FG}}$ captures the main structure of $G_{k}$, while the approximation error $\|G_{k}-P_{\mathcal{S}_{k}^{\text{FG}}}(G_{k})\|$ is governed by the spectral tail of $G_{k}$. Crucially, this tail is independent of the rank gap in the objective, underscoring a key distinction between representation deficiency (e.g., LoRA) and update efficiency. In practice, given $(U_{k},V_{k})$ from the rank-$r$ SVD of $G_{k}$, one can also restart with $\mathcal{M}_{k}(\xi_{k}^{\text{new}})=-\eta_{k}U_{k}V_{k}$, which corresponds to the update from the recent training benchmark Muon (Jordan et al., 2024), providing improved stability over (8).
[…]
To implement Algorithm 2, directly assigning $(A_{k-1},B_{k-1})\leftarrow 1/{\sqrt{\gamma}}(U_{k}\sqrt{\Lambda_{k}},\sqrt{\Lambda_{k}}V_{k})$ can cause instability due to mismatches in optimization states. A similar instability issue has also been reported for GaLore (Chen et al., 2024). For robustness, we propose alignment techniques to maintain consistency of subspace bases, momentum, and velocity. In particular, restarts often produce gradients with much larger magnitudes, leaving the Adam velocity “too cold” and causing unstable steps. We therefore apply a lightweight velocity alignment, rescaling the velocity to match the new gradient magnitude, $v\leftarrow\|g\|^{2}/\|v\|\,v$, together with a short $\beta_{2}$ warm-up, which we find to be the most important component for stable performance in practice; see more details in Appendix D. Finally, we remark both empirical evidence and theoretical results suggest that gradients $G_{k}$ in deep learning often have strong low-rank structure, making them especially suitable for efficient SVD-based approximations (Cosson et al., 2023; Yang et al., 2023; Zhao et al., 2024).
3.2 Exploitation via SVD Subspace
Having discussed exploration techniques inspired by subspace minimization, we now turn to the complementary philosophy: exploitation within the current subspace.
Algorithm 3 PESO-LoRA-T: PESO with LoRA and Subspace ExploiTation
1:Pretrained weights $W_{0}\in\mathbb{R}^{m\times n}$, initial subspace matrices $U_{0}\in\mathbb{R}^{m\times r}$, $V_{0}\in\mathbb{R}^{r\times n}$, initial coordinate $\xi_{0}\in\mathbb{R}^{r}$, frequency $K$.
2:Set $k\leftarrow 1$.
3:while stopping criterion not met do
4:  Keep $(U_{k},V_{k})\leftarrow(U_{k-1},V_{k-1})$.
5:  if $k-1\bmod K=0$ then
6:   $(U_{k},V_{k})\leftarrow\texttt{AdamW}(U_{k-1},V_{k-1})$.
7:  end if
8:  $\xi_{k}\leftarrow\texttt{AdamW}(\xi_{k-1})$.
9:  $k\leftarrow k+1$.
10:end while
11:return $W_{0}+U_{k}\operatorname{diag}(\xi_{k})V_{k}$.
As outlined in Section 2.2, an SVD-based parameterization provides a clean and principled way to define $\mathcal{M}_{k}$. Specifically, we approximate the target mapping $\mathcal{M}(\xi^{*})$ by a sum of rank-$r$ components, $\sum_{k}U_{k}\xi^{*}_{k}V_{k}$, where each pair $(U_{k},V_{k})$ defines an SVD subspace of rank $r$. Because SVD naturally captures the dominant gradient directions, this parameterization ensures that exploitation is focused on the most informative directions in the weight space.
Within each subspace, we optimize the low-dimensional $\xi$ for $K$ steps using AdamW. This design can be viewed as an extension of LoRA, with the key difference that the SVD structure explicitly decouples subspace exploitation (through $\xi$) from exploration (through $(U,V)$). The practical variant is summarized in Algorithm 3. A small frequency $K$ (e.g., $1$ or $2$) often suffices for strong performance without significant overhead. A detailed version of Algorithm 3 can be found in Appendix E.
4 Experiments
In this section, we conduct experiments to evaluate our methods across diverse tasks and models, comparing with standard LoRA-based approaches and full fine-tuning. We first assess natural language understanding on the GLUE benchmark (Wang et al., 2018) by fine-tuning T5-base (Raffel et al., 2020). We then evaluate natural language generation on Llama models for tasks including mathematical reasoning, code generation, and general instruction tuning. Finally, we demonstrate that LoRA-PESO-R remains effective even under strict memory constraints when trained for more epochs. Implementation details are provided in Appendix F.
Table 2: Fine-tuned T5-base on natural language understanding tasks with rank set to 8. Results are reported as accuracy (%) over 3 runs. Bold and underline indicate the highest and second-highest accuracies excluding PESO-LoRA-T, which is shaded in gray and omitted from direct comparison due to its longer runtime.
Method MNLI SST-2 CoLA QNLI MRPC
LoRA 85.30±0.04 94.04±0.09 72.84±1.25 93.02±0.07 68.38±0.01
LoRA+ 85.81±0.09 93.85±0.24 77.53±0.20 93.14±0.03 74.43±1.39
P-LoRA 85.28±0.15 93.88±0.11 79.58±0.67 93.00±0.07 83.91±1.16
PiSSA 85.75±0.07 94.07±0.06 74.27±0.39 93.15±0.14 76.31±0.51
LoRA-GA 85.70±0.09 94.11±0.18 80.57±0.20 93.18±0.06 85.29±0.24
LoRA-Pro 86.03±0.19 94.19±0.13 81.94±0.24 93.42±0.05 86.60±0.14
LoRA-One 85.89±0.08 94.53±0.13 82.04±0.22 93.37±0.02 87.83±0.37
PESO-LoRA-R 86.08±0.15 94.61±0.09 81.50±0.16 93.43±0.06 86.36±0.11
[…]
4.1 Natural Language Understanding Tasks
We fine-tune the T5-base model on a subset of GLUE, including MNLI, SST-2, CoLA, QNLI, and MRPC, and evaluate performance using test accuracy (%). Following the setting in (Zhang et al., 2025b), we compare our method against several LoRA variants, including LoRA (Hu et al., 2022), LoRA+ (Hayou et al., 2024), P-LoRA (Zhang and Pilanci, 2024), PiSSA (Meng et al., 2024), LoRA-GA (Wang et al., 2024a), LoRA-Pro (Wang et al., 2024b), and LoRA-One (Zhang et al., 2025b). For fairness, hyperparameters are tuned individually for each method.
The results are summarized in Table 2. PESO-LoRA-R and PESO-LoRA-T achieve the best performance on three of the five tasks, which also have larger datasets. On the remaining tasks, PESO-LoRA-T ranks second. This demonstrates the overall efficiency and robustness of our approaches, with advantages most evident on larger datasets that demand longer training and stronger exploration–exploitation. Moreover, PESO-LoRA-T generally outperforms PESO-LoRA-R, but at the cost of $1.4\times$ more computation time, whereas PESO-LoRA-R runs at nearly the same speed as standard LoRA. Memory costs are comparable across all methods, so the choice ultimately depends on whether performance or efficiency is prioritized.
4.2 Natural Language Generation Tasks
We evaluate our method on the more capable Llama-3.1-8B (Dubey et al., 2024) model across three downstream tasks. For mathematical reasoning, we fine-tune on a 100k subset of MetaMathQA (Yu et al., 2023) and evaluate on GSM8K (Cobbe et al., 2021). For general instruction tuning, we fine-tune on Alpaca (Taori et al., 2023) and evaluate on MMLU (Hendrycks et al., 2020). For code generation, we use a 100k subset of Code-Feedback (Zheng et al., 2024) and evaluate on HumanEval (Chen et al., 2021), reporting PASS@1. To ensure fairness, all datasets are preprocessed to exclude overlaps with test sets.
The results are reported in Table 3. PESO-LoRA-R consistently outperforms all LoRA variants across all benchmarks on this advanced model. Notably, the gains are most pronounced in mathematical reasoning and code generation—tasks which involves larger fine-tuning datasets, highlighting the substantial advantages of subspace exploration and exploitation for complex tasks. For completeness, Table 7 in Appendix F.2 reports results on Llama-2-7B (Touvron et al., 2023), evaluating both PESO-LoRA-T and PESO-LoRA-R, where our methods similarly achieve competitive performance against established baselines. We further compare against pretraining-oriented approaches, including GaLore (Zhao et al., 2024), Fira (Chen et al., 2024), and APOLLO (Zhu et al., 2024); Table 5 shows that our PEFT-specialized PESO-LoRA-R consistently outperforms these pretraining baselines in the fine-tuning regime.
We conduct ablations over the rank $r$ and restart frequency $K$. Results in Appendix F.3 show that PESO-LoRA-R achieves strong performance with small ranks and remains robust across a wide range of $K$. Appendix F.3 also reports memory and runtime comparisons, showing a memory footprint similar to vanilla LoRA and comparable overall computational cost. In particular, Table 11 shows that the additional SVD cost over LoRA is only 1.9% of total runtime in our experiments, indicating negligible overhead.
Table 3: Fine-tuned Llama-3.1-8B on natural language generation tasks with rank set to 8. Results are reported as accuracy (%) over 3 runs. Bold and underline indicate the highest and second-highest accuracies.
LoRA LoRA-GA LoRA-One PESO-LoRA-R
GSM8K 70.64±0.53 76.67±0.31 77.71±0.17 77.79±0.18
MMLU 63.95±0.05 62.91±0.08 64.33±0.14 64.34±0.21
HumanEval 42.47±2.56 44.32±5.64 45.32±1.52 47.15±0.76
Table 4: Fine-tuned T5-base (4 epochs) on the SST-2 dataset. Results are reported as accuracy (%) over 3 runs. Bold and underline indicate the highest and second-highest accuracies.
Method Epoch 1 Epoch 2 Epoch 3 Epoch 4
LoRA ($r=2$) $93.02_{\pm 0.44}$ $93.47_{\pm 0.51}$ $93.41_{\pm 0.12}$ $93.52_{\pm 0.17}$
LoRA ($r=4$) $94.23_{\pm 0.30}$ $94.42_{\pm 0.15}$ $94.46_{\pm 0.10}$ $94.61_{\pm 0.25}$
LoRA ($r=8$) $93.85_{\pm 0.30}$ $94.03_{\pm 0.09}$ $94.54_{\pm 0.05}$ $94.54_{\pm 0.23}$
PESO-LoRA-R ($r=2$) $\underline{94.30}_{\pm 0.15}$ $\underline{94.47}_{\pm 0.08}$ $\underline{94.84}_{\pm 0.25}$ $\textbf{95.14}_{\pm 0.15}$
Full fine-tuning $\textbf{94.42}_{\pm 0.11}$ $\textbf{94.70}_{\pm 0.10}$ $\textbf{94.85}_{\pm 0.11}$ $\underline{94.90}_{\pm 0.06}$
Table 5: Comparison with pretraining-oriented subspace methods. Fine-tuned Llama-3-8B-Instruct (Alpaca-en-demo, $r=8$). Results are MMLU accuracy (%). Baseline results from Zhu et al. (2024). Bold and underline indicate the highest and second-highest accuracies.
Method STEM Social Sciences Humanities Others Average
GaLore 54.50 75.11 58.59 72.03 64.43
Fira 53.53 75.46 58.59 72.09 64.32
APOLLO /w SVD 54.73 75.46 58.72 72.68 64.76
APOLLO 54.37 75.86 58.18 71.69 64.35
PESO-LoRA-R 57.15 76.86 60.79 73.57 66.47
4.3 Multi-Epoch Low-Rank Analysis
To highlight the effect of subspace exploration, we extend T5-base fine-tuning on SST-2 from one epoch (Section 4.1) to four. The longer schedule allows more thorough exploration and mitigates the intrinsic low-rank bottleneck. As shown in Table 4, PESO-LoRA-R with $r=2$ consistently outperforms standard LoRA even with larger ranks ($r=4,8$), indicating that it alleviates the low-rank limitation and achieves stronger performance under tightly restricted memory budgets.
5 Convergence Analysis
In this section we establish convergence guarantees for PESO-LoRA-R, a practical instantiation of PESO. A more general asymptotic convergence result for PESO is provided in Appendix G. We consider a nonconvex objective $\ell(W)$ and make the following standard assumptions.
Assumption 1.
The loss $\ell$ is bounded below and has $L$-Lipschitz continuous gradients.
Assumption 2.
Stochastic gradients $\widetilde{G}_{k}$ of the full gradient $G_{k}=\nabla\ell(W_{k})$ are accessible and satisfy $\mathbb{E}(\widetilde{G}_{k})=G_{k}$. There exists $C>0$ such that $\mathbb{V}(\widetilde{G}_{k})\leq C$.
We analyze PESO-LoRA-R (Algorithm 2), where the inner optimizer is set as Adam. Below we state a standard regularity assumption on the low-rank objective $f(A,B)=\ell(W+AB)$ optimized by Adam.
Assumption 3.
At each iteration, PESO-LoRA-R applies Adam to the low-rank objective $f(A,B)=\ell(W+AB)$. We assume for $f$: (i) $\nabla f$ is Lipschitz continuous, (ii) stochastic gradients are unbiased, and (iii) stochastic gradient variance is bounded.
Let $P_{r}(G_{k})$ denote the rank-$r$ truncated SVD of $G_{k}$. Since $P_{r}(\cdot)$ is nonlinear, in general $\mathbb{E}[P_{r}(\widetilde{G}_{k})]\neq P_{r}(\mathbb{E}[\widetilde{G}_{k}])$. Our analysis only requires that this projection-induced bias is asymptotically negligible in a stepsize-weighted sense.
Assumption 4.
At each restart step $(k-1)\bmod K=0$, the estimated subspace satisfies $\|\mathbb{E}[P_{r}(\widetilde{G}_{k})]-P_{r}(G_{k})\|_{F}\leq\varepsilon_{k},\sum_{k}\eta_{k}\varepsilon_{k}^{2}<\infty.$
Remark 1.
This assumption holds in several standard regimes in subspace optimization literature. It holds trivially in the full-batch setting ($\varepsilon_{k}=0$) and is satisfied in the large-batch regime via concentration bounds $\varepsilon_{k}=\mathcal{O}(|\mathcal{B}|^{-1/2})$; see, e.g., (He et al., 2024). Furthermore, using independent randomized subspaces (Chen et al., 2025) yields $\varepsilon_{k}=0$ due to conditional unbiasedness.
Theorem 5.1.
Under Assumptions 14 and following the hyperparameter choices from Appendix H, the iterates generated by PESO-LoRA-R satisfy
$\min_{0\leq k\leq T-1}\mathbb{E}\|\nabla\ell(W_{k})\|_{2}^{2}=\mathcal{O}\!\left(1/{\sqrt{T}}\right).$
6 Conclusions and Limitations
This paper bridges classical methodology from nonlinear optimization with the practical challenge of memory-efficient LLM fine-tuning. It highlights two key perspectives: 1) practical constraints in LLM training, such as memory limits, can motivate specialized optimization designs; 2) principles from nonlinear optimization can in turn guide the development of practical algorithms for LLMs. We believe this opens promising directions for principled and scalable LLM training, while underscoring a broader philosophy: the rapid progress in LLMs can be enriched by classical foundations in computation and optimization. Our study has certain limitations. Due to limited resources, our experiments are restricted to medium-scale settings and do not yet reach the largest practical regimes. Extending our framework to full-scale pre-training remains an important future work, and we expect the methodology developed here to provide a solid foundation for such efforts.
Acknowledgments
We would like to thank Professor Jorge Nocedal for his invaluable insights and discussions, which greatly assisted in formulating the framework and refining the theoretical development of this work.
Appendix A Synthetic Example of LoRA Deficiency
One critical limitation in the literature is the absence of convergence guarantees toward valid optimality conditions of (1). Most existing works establish convergence only with respect to the low-dimensional parameters—such as the factors $A$ and $B$ in LoRA—but do not address convergence with respect to the full-parameter matrix $W$. For instance, Jiang et al. (2024) shows that $\nabla_{A}\ell(W_{0}+A_{k}B_{k})$ and $\nabla_{B}\ell(W_{0}+A_{k}B_{k})$ vanish as $k\to\infty$, while leaving the behavior of $\nabla_{W}\ell$ uncharacterized. This gap is not merely technical: it highlights a fundamental deficiency of PEFT methods compared to standard full-parameter training. In fact, the optimal loss attained by LoRA can be arbitrarily worse than the true optimal loss of (1). To illustrate this, consider the following simple synthetic example in matrix optimization:
$\min_{W\in\mathbb{R}^{n\times n}}\quad\|W-M\|_{F}^{2},\quad\text{ where }M=a\cdot\operatorname{diag}(1,\dots,1,0,\dots,0),\quad(r{+}1\text{ ones}).$ (9)
The optimal solution is clearly $W^{*}=M$ with $f(W^{*})=0$. However, applying LoRA with rank $r$ to (9) can at most achieve a rank-$r$ approximation of $M$, and attains $f(A^{*}B^{*})=a^{2}$. As $a$ increases, or as the rank mismatch between the LoRA adapters and the true solution grows, the optimality gap between LoRA and full gradient methods can become arbitrarily large.
This example underscores the cost of memory restrictions: while low-rank parameterizations save memory, they may fundamentally limit convergence to the true optimum. In the middle and right panels of Figure 1, we show that LoRA with exploration (PESO-LoRA-R) can effectively converge to the true optimal solution while LoRA would not. Note in this implementation of PESO-LoRA-R (Algorithm 2), the SVD of the full gradient is a rank-$r$ SVD and therefore the low-rankness of this computational scheme would not affect the convergence.
Appendix B Review on Subspace Minimization
It is worth noting that Algorithm 1 is essentially equivalent to the classical two-loop subspace minimization scheme (Conn et al., 1994), summarized in Algorithm 4.
The key distinction between warm-start and restart, discussed in (5) and (6) of Section 2.1, lies in how the subproblem (10) is initialized within each outer iteration of Algorithm 4.
Algorithm 4 Classical Iterated-Subspace Minimization
1:Initialization $W_{0}\in\mathbb{R}^{m\times n}$, $\xi_{0}\in\mathbb{R}^{d}$, $\mathcal{M}_{0}$, an algorithmic subroutine UpdateM, an optimizer Opt, frequency $K$.
2:Set $k\leftarrow 1$ and $\widetilde{W}_{0}\leftarrow W_{0}$.
3:while stopping criteria not satisfied do
4:  Update the subspace $(\mathcal{M}_{k},\widetilde{W}_{k})\leftarrow\texttt{UpdateM}(\mathcal{M}_{k-1},\widetilde{W}_{k-1})$.
5:  Approximately solve the subspace minimization by Opt using $K$ inner-loop iterations:
$\xi_{k}^{*}\leftarrow\text{approx}\arg\min_{\xi}\,\,\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi))$ (10)
6:  $W_{k}\leftarrow\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k}^{*})$.
7:  $k\leftarrow k+1$.
8:end while
Appendix C Projected Subspace and Memory Efficiency
One-sided projected subspaces in PESO can offer stronger memory efficiency than LoRA. This idea is exemplified by GaLore (Zhao et al., 2024), which we now place in the PESO framework. GaLore requires memory of order $mn+mr+2nr$ (assuming $m\leq n$), compared to LoRA’s $mn+3mr+3nr$.
With the projected subspace representation in Table 1, optimization reduces to $\xi:=R\in\mathbb{R}^{r\times n}$. By the chain rule,
$\nabla_{R}\ell(W_{k})=\nabla_{\xi}\ell(\widetilde{W}_{k}+P_{k}R_{k})=P_{k}^{\top}\nabla_{W}\ell(W_{k})=P_{k}^{\top}G_{k}.$ (11)
Once $P_{k}$ is computed and stored, subspace gradients are obtained directly from $G_{k}$ with no extra overhead, though computing the full $G_{k}$ each iteration is more costly than subspace-only gradients.
Then, an exploitation step in the subspace by gradient descent with learning rate $\eta_{k}$ gives
$W_{k+1}$ $=\widetilde{W}_{k}+\mathcal{M}_{k}(R_{k}-\eta_{k}P_{k}^{\top}G_{k})$ (12)
$=\widetilde{W}_{k}+P_{k}(R_{k}-\eta_{k}P_{k}^{\top}G_{k})$
$=W_{k}-\eta_{k}P_{k}P_{k}^{\top}G_{k},$
which matches the classical projected subspace descent step (Kozak et al., 2019).
In GaLore, $P_{k}$ is chosen as the rank-$r$ left SVD of the full gradient at fixed intervals. PESO recovers GaLore when the subspace gradient $\nabla_{R}\ell(W_{k})=P_{k}^{\top}G_{k}$ in (12) is replaced with its Adam update.
This shows how GaLore saves memory: gradients of $\xi_{k}$ are derived directly from $G_{k}$, and updates are written back into $W$ via (12), reusing the stored pretrained weights $W_{0}$. Thus explicit storage of $\xi_{k}$ is unnecessary. Further savings arise because GaLore omits optimizer states for $\mathcal{M}_{k}$ (i.e., for $P_{k}$), instead updating $P_{k}$ by direct reassignment in a restart manner. However, smoother transitions of subspace parameters often yield greater stability, as observed in our experiments; we discuss smoothing techniques for restart strategies in the next section.
Appendix D Implementation of PESO-LoRA-R
This section details implementation strategies for PESO-LoRA-R that are critical for practical stability, specifically regarding subspace smoothing and the alignment of optimizer states. While these techniques collectively enhance robustness, our empirical observations indicate that velocity alignment is the most significant factor in stabilizing training dynamics during restarts. Consequently, for a streamlined implementation that faithfully recovers our results, we strongly recommend prioritizing the velocity alignment mechanism described in Section D.2.
D.1 Smoothing the Subspaces
A potential issue with restart methods (see (6)) is that they directly reassign the subspace parameterization from new information, which can introduce sharp changes and instability, especially in LLM training where stochastic noise is significant.
To mitigate this, we adopt an Exponential Moving Average (EMA) of old and new subspaces, similar in spirit to how Adam (Kingma and Ba, 2014) stabilizes noisy updates. However, this is nontrivial in PESO-LoRA-R (Algorithm 2), since the pre-restart adapters $(A_{k},B_{k})$—evolved through Adam dynamics—may differ significantly in scale and coordinates from the restarted pair $(U_{k}\sqrt{\Lambda_{k}},\sqrt{\Lambda_{k}}V_{k})$ obtained from rank-$r$ SVD. A naive EMA would mismatch these terms and discard valuable exploration information.
We resolve this by performing basis and scaling alignment. For clarity, we omit the subscript $k$. Given current adapters $(A,B)$, we first compute thin QR factorizations:
$A=Q_{A}R_{A},\quad B=R_{B}Q_{B}^{\top},\quad Q_{A}^{\top}Q_{A}=I_{r},\;Q_{B}^{\top}Q_{B}=I_{r},$
to extract bases $(Q_{A},Q_{B})$ and decouple scaling. Next, let the rank-$r$ SVD of the full gradient be $G\approx U\,\Sigma\,V^{\top}$. We align $(U,V)$ to $(Q_{A},Q_{B})$ by applying SVD to $Q_{A}^{\top}U$ and $Q_{B}^{\top}V$:
$Q_{A}^{\top}U=P_{U}\Sigma_{U}Q_{U}^{\top},\;\;R_{U}:=P_{U}Q_{U}^{\top},\qquad Q_{B}^{\top}V=P_{V}\Sigma_{V}Q_{V}^{\top},\;\;R_{V}:=P_{V}Q_{V}^{\top},$
yielding aligned bases
$\widehat{U}:=UR_{U}^{\top},\qquad\widehat{V}:=VR_{V}^{\top}.$
Here $R_{U}$ solves
$\min_{R}\|UR-Q_{A}\|_{F},\quad RR^{\top}=I,$
so $\widehat{U}$ best aligns $U$ with $Q_{A}$ in Frobenius norm; the same holds for $\widehat{V}$. This produces the best alignment by classical low-rank SVD guarantees.
We then smooth the bases via EMA:
$U_{\mathrm{ema}}:=\tau_{1}Q_{A}+(1-\tau_{1})\widehat{U},\qquad V_{\mathrm{ema}}:=\tau_{1}Q_{B}+(1-\tau_{1})\widehat{V}$,
for smoothing parameter $\tau_{1}$. To smooth the scaling, we project the old adapter into the new bases and combine with the gradient:
$S_{\mathrm{new}}:=\tau_{2}\,[U_{\mathrm{ema}}^{\top}(AB)\,V_{\mathrm{ema}}]-(1-\tau_{2})\,[U_{\mathrm{ema}}^{\top}G\,V_{\mathrm{ema}}],$
with parameter $\tau_{2}$. Since this merges the scaling of $A$ and $B$, we refactorize $S_{\mathrm{new}}$ using polar decomposition (Glentis et al., 2025):
$S_{\mathrm{new}}=R_{L}\,\Sigma\,R_{R}^{\top},$
and define the new adapters via balanced splitting:
$A_{\mathrm{new}}:=U_{\mathrm{ema}}\,R_{L}\,\Sigma^{1/2},\qquad B_{\mathrm{new}}:=\Sigma^{1/2}\,R_{R}^{\top}\,V_{\mathrm{ema}}^{\top}.$
Finally, although empirically less effective, we also rotate the momentum vectors according to the new basis. Specifically, we compute
$T_{A}:=Q_{A}^{\top}U_{\mathrm{ema}},\qquad T_{B}:=Q_{B}^{\top}V_{\mathrm{ema}},$
and update the pre-restart momentum $(m_{A},m_{B})$ as
$m_{A}\leftarrow m_{A}T_{A},\qquad m_{B}\leftarrow T_{B}^{\top}m_{B}.$
Training then proceeds with the new adapters $(A_{\mathrm{new}},B_{\mathrm{new}})$ in PESO-LoRA-R.
D.2 Momentum and Velocity Alignment
Even with basis and scaling alignment from the previous subsection, another stability issue arises: the new adapters $(A_{\mathrm{new}},B_{\mathrm{new}})$ can induce gradients of very different magnitudes compared to the old $(A,B)$. Since restarts are based on the SVD of the full gradient, the new adapters align with top gradient directions, so the gradients with respect to $(A_{\mathrm{new}},B_{\mathrm{new}})$ are typically larger. This mismatch can leave the velocity “too cold”: historical states $(v_{A},v_{B})$ may underestimate the new gradient magnitudes, leading to an excessively large normalized step and unstable behavior, often observed as jumps in the loss curve.
To address this, we propose a combined momentum/velocity scaling technique with a $\beta_{2}$ warm-up. Let $(m_{A},m_{B})$ and $(v_{A},v_{B})$ denote the momentum and velocity before restart, and $(g_{A},g_{B})$ the gradients after restart computed with respect to $(A_{\mathrm{new}},B_{\mathrm{new}})$. We define scaling factors
$s^{(v)}_{A}\;=\;\frac{\|g_{A}\|^{2}}{{\|v_{A}\|}},\quad s^{(m)}_{A}\;=\;\frac{\|g_{A}\|}{\|m_{A}\|},\quad s^{(v)}_{B}\;=\;\frac{\|g_{B}\|^{2}}{{\|v_{B}\|}},\quad s^{(m)}_{B}\;=\;\frac{\|g_{B}\|}{\|m_{B}\|},$
which correct scale mismatches between the old optimization states and the new gradients. Here, $\|\cdot\|$ denotes the RMS norm. Momentum and velocity are then rescaled as
$v_{A}\;\leftarrow\;s_{A}^{(v)}\,v_{A},\quad m_{A}\;\leftarrow\;s_{A}^{(m)}\,m_{A},\quad v_{B}\;\leftarrow\;s_{B}^{(v)}\,v_{B},\quad m_{B}\;\leftarrow\;s_{B}^{(m)}\,m_{B}.$
This resolves scale mismatches, but an additional adjustment is needed: $\beta_{2}=0.999$ (velocity EMA) adapts much more slowly than $\beta_{1}=0.9$ (momentum EMA). At initialization, bias correction balances these, but after a restart we require extra correction. We therefore decrease $\beta_{2}$ immediately after a restart and gradually warm it back to $0.999$ over a window $T$. If a restart occurs at iteration $t_{r}$, then for $t_{r}\leq t\leq t_{r}+T$ we set
$\beta_{2}(t)\;=\;\beta_{2,\min}+\big(\beta_{2,\text{final}}-\beta_{2,\min}\big)\,\tfrac{1}{2}\!\left(1-\cos\tfrac{\pi(t-t_{r})}{T}\right),\quad\beta_{2,\text{final}}=0.999,$
and for $t>t_{r}+T$ we use $\beta_{2}=0.999$ as usual. In our experiments, we set $\beta_{2,\min}=0.95$ and $T=\lfloor K/3\rfloor$.
Appendix E Detailed Version of Algorithms
In this section, we provide detailed versions of PESO-LoRA-R (Algorithm 2) and PESO-LoRA-T (Algorithm 3) which are implemented in the experiments of Section 4.
Algorithm 5 PESO-LoRA-R: PESO with LoRA and Subspace ExploRation (Practical Version)
1:Pre-trained parameters $W_{0}\in\mathbb{R}^{m\times n}$, frequency $K$, scale parameter $\gamma$, learning rate $\eta$, AdamW hyperparameters $(\beta_{1},\beta_{2},\epsilon,\lambda)$.
2:Set $k\leftarrow 1$, $\widetilde{W}_{0}\leftarrow W_{0}$, $A_{0}\leftarrow 0$, $B_{0}\leftarrow 0$.
3:Initialize AdamW states $m_{A}\leftarrow 0$, $v_{A}\leftarrow 0$, $m_{B}\leftarrow 0$, $v_{B}\leftarrow 0$.
4:while stopping criteria not satisfied do
5:  if $k-1\mod{K}=0$ then
6:   $\widetilde{W}_{k}\leftarrow\widetilde{W}_{k-1}+A_{k-1}B_{k-1}$.
7:   Compute stochastic full gradient $G_{k}$.
8:   $(U_{k},\Lambda_{k},V_{k})\leftarrow{\tt SVD}(-G_{k})$. $\triangleright$ Top-$r$ SVD of $G_{k}$
9:   Set $(A_{k-1},B_{k-1})\leftarrow(A_{\text{new}},B_{\text{new}})$ according to Appendix D.1.
10:   Update $m_{A},m_{B},v_{A},v_{B},\beta_{2}$ according to Appendix D.2.
11:  end if
12:  Compute gradients $g_{A,k}\leftarrow\nabla_{A}\ell(\widetilde{W}_{k}+A_{k-1}B_{k-1})$, $g_{B,k}\leftarrow\nabla_{B}\ell(\widetilde{W}_{k}+A_{k-1}B_{k-1})$.
13:  $m_{A}\leftarrow\beta_{1}m_{A}+(1-\beta_{1})g_{A,k}$, $v_{A}\leftarrow\beta_{2}v_{A}+(1-\beta_{2})g_{A,k}\odot g_{A,k}$.
14:  $m_{B}\leftarrow\beta_{1}m_{B}+(1-\beta_{1})g_{B,k}$, $v_{B}\leftarrow\beta_{2}v_{B}+(1-\beta_{2})g_{B,k}\odot g_{B,k}$.
15:  $\widehat{m}_{A}\leftarrow m_{A}/(1-\beta_{1}^{k})$, $\widehat{v}_{A}\leftarrow v_{A}/(1-\beta_{2}^{k})$.
16:  $\widehat{m}_{B}\leftarrow m_{B}/(1-\beta_{1}^{k})$, $\widehat{v}_{B}\leftarrow v_{B}/(1-\beta_{2}^{k})$.
17:  $A_{k}\leftarrow A_{k-1}-\eta\lambda A_{k-1}-\eta\,\widehat{m}_{A}/(\sqrt{\widehat{v}_{A}}+\epsilon)$.
18:  $B_{k}\leftarrow B_{k-1}-\eta\lambda B_{k-1}-\eta\,\widehat{m}_{B}/(\sqrt{\widehat{v}_{B}}+\epsilon)$.
19:  $k\leftarrow k+1$.
20:end while
21:return $\widetilde{W}_{k}+A_{k}B_{k}$.
Algorithm 6 PESO-LoRA-T: PESO with LoRA and Subspace ExploiTation (Practical Version)
1:Pretrained weights $W_{0}\in\mathbb{R}^{m\times n}$, initial subspace matrices $U_{0}\in\mathbb{R}^{m\times r}$, $V_{0}\in\mathbb{R}^{r\times n}$, initial coordinate $\xi_{0}\in\mathbb{R}^{r}$, frequency $K$, learning rate $\eta$, AdamW hyperparameters $(\beta_{1},\beta_{2},\epsilon,\lambda)$.
2:Set $k\leftarrow 1$.
3:Initialize AdamW states $m_{U},m_{V},m_{\xi}\leftarrow 0$ and $v_{U},v_{V},v_{\xi}\leftarrow 0$.
4:while stopping criterion not met do
5:  Keep $(U_{k},V_{k})\leftarrow(U_{k-1},V_{k-1})$.
6:  if $k-1\bmod K=0$ then
7:   Compute gradient $g_{U,k}\leftarrow\nabla_{U}\ell(W_{0}+U_{k-1}\operatorname{diag}(\xi_{k-1})V_{k-1})$.
8:   Compute gradient $g_{V,k}\leftarrow\nabla_{V}\ell(W_{0}+U_{k-1}\operatorname{diag}(\xi_{k-1})V_{k-1})$.
9:   $m_{U}\leftarrow\beta_{1}m_{U}+(1-\beta_{1})g_{U,k}$, $v_{U}\leftarrow\beta_{2}v_{U}+(1-\beta_{2})g_{U,k}\odot g_{U,k}$.
10:   $m_{V}\leftarrow\beta_{1}m_{V}+(1-\beta_{1})g_{V,k}$, $v_{V}\leftarrow\beta_{2}v_{V}+(1-\beta_{2})g_{V,k}\odot g_{V,k}$.
11:   $\widehat{m}_{U}\leftarrow m_{U}/(1-\beta_{1}^{k})$, $\widehat{v}_{U}\leftarrow v_{U}/(1-\beta_{2}^{k})$.
12:   $\widehat{m}_{V}\leftarrow m_{V}/(1-\beta_{1}^{k})$, $\widehat{v}_{V}\leftarrow v_{V}/(1-\beta_{2}^{k})$.
13:   $U_{k}\leftarrow U_{k-1}-\eta\lambda U_{k-1}-\eta\,\widehat{m}_{U}/(\sqrt{\widehat{v}_{U}}+\epsilon)$.
14:   $V_{k}\leftarrow V_{k-1}-\eta\lambda V_{k-1}-\eta\,\widehat{m}_{V}/(\sqrt{\widehat{v}_{V}}+\epsilon)$.
15:  end if
16:  Compute gradient $g_{\xi,k}\leftarrow\nabla_{\xi}\ell(W_{0}+U_{k}\operatorname{diag}(\xi_{k-1})V_{k})$.
17:  $m_{\xi}\leftarrow\beta_{1}m_{\xi}+(1-\beta_{1})g_{\xi,k}$, $v_{\xi}\leftarrow\beta_{2}v_{\xi}+(1-\beta_{2})g_{\xi,k}\odot g_{\xi,k}$.
18:  $\widehat{m}_{\xi}\leftarrow m_{\xi}/(1-\beta_{1}^{k})$, $\widehat{v}_{\xi}\leftarrow v_{\xi}/(1-\beta_{2}^{k})$.
19:  $\xi_{k}\leftarrow\xi_{k-1}-\eta\lambda\xi_{k-1}-\eta\,\widehat{m}_{\xi}/(\sqrt{\widehat{v}_{\xi}}+\epsilon)$.
20:  $k\leftarrow k+1$.
21:end while
22:return $W_{0}+U_{k}\operatorname{diag}(\xi_{k})V_{k}$.
Appendix F Experimental Details
All experiments are conducted on NVIDIA RTX A6000 GPUs. For PESO-LoRA-R, to further reduce computational cost, we restrict the exploration frequency to at most two times per epoch.
F.1 Natural Language Understanding: Hyperparameter Settings
In Section 4.1, we present the results of our methods and various LoRA-based algorithms on natural language understanding tasks, following the prompt tuning configuration of (Wang et al., 2024a). The general hyperparameter settings are kept consistent across all algorithms which are shown in Table 6. To ensure a fair comparison, we follow (Zhang et al., 2025b) and tune the learning rates via grid search over $\{1\times 10^{-4},2\times 10^{-4},5\times 10^{-4},1\times 10^{-3}\}$. Additionally, following the choices of Zhang et al. (2025b), the scale parameters for LoRA-One are set to be {128, 16, 128, 128, 64} for {MNLI, SST-2, CoLA, QNLI, MRPC}.
For PESO-LoRA-R, we set the smoothing parameter $\tau_{1}=\tau_{2}=0.9$, with frequency $K$ chosen as $\{2000,500,100,500,40\}$ for {MNLI, SST-2, CoLA, QNLI, and MRPC} based on empirical observations. When $(k-1)\bmod K=0$ and $k\neq 0$, we set the scale parameter $\gamma=1$; when $k=0$, the scale parameter is set the same as in LoRA-One. To further reduce computational cost, we restrict the restart frequency to times per epoch. For PESO-LoRA-T, we set frequency $K=1$ for all datasets.
Table 6: Common hyperparameters for LoRA fine-tuning on T5-base model.
Epoch Optimizer $(\beta_{1},\beta_{2})$ $\epsilon$ Batch Size Weight Decay LR Scheduler
1 AdamW (0.9, 0.999) $1\times 10^{-8}$ 32 0 cosine
Warm-up Ratio LoRA Alpha #Runs Sequence Length Adapt Precision Backbone Precision Gradient Batch Size
0.03 16 3 128 FP32 FP32 8
F.2 Natural Language Generation: Hyperparameter Settings
Table 7: Performance of fine-tuned Llama-2-7B on natural language generation tasks with rank set to 8. Results are reported as accuracy (%) over 3 runs. Bold and underline indicate the highest and second-highest accuracies excluding PESO-LoRA-T.
[…]
F.2.1 Comparison with LoRA-based methods
For natural language generation tasks in Section 4.2, we follow the configuration of prompt tuning and strategy of hyperparameter tuning in (Zhang et al., 2025b) to ensure fair comparison. We search the best learning rate over $\{5\times 10^{-4},\,2\times 10^{-4},\,1\times 10^{-4},\,5\times 10^{-5},\,2\times 10^{-5},\,1\times 10^{-5}\}$, and the general hyperparameter setting is summarized in Table 8. Additionally, following the choice of Zhang et al. (2025b), the scale parameters are tuned within $\{16,32,64,128\}$ for LoRA-GA, LoRA-One and our methods to achieve the best performances.
Table 8: Common hyperparameters for LoRA fine-tuning on Llama-2-7B and LLama-3.1-8B model.
Epoch Optimizer $(\beta_{1},\beta_{2})$ $\epsilon$ Batch Size Weight Decay LR Scheduler
1 AdamW (0.9, 0.999) $1\times 10^{-8}$ 32 0 cosine
Warm-up Ratio LoRA Alpha #Runs Sequence Length Adapter Precision Backbone Precision Gradient Batch Size
0.03 16 3 1024 FP32 BF16 8
For PESO-LoRA-R, we set the smoothing parameter $\tau_{1}=\tau_{2}=0.9$, with frequency $K=500$ for all the experiments. When $(k-1)\bmod K=0$ and $k\neq 0$, we set the scale parameter $\gamma=1$; when $k=0$, the scale parameter is set the same as in LoRA-One. For PESO-LoRA-T, we set frequency $K=1$ for all datasets.
F.2.2 Comparison with pretraining-oriented subspace methods
[…]
The results are shown in Table 5, demonstrating that our PEFT-specialized method PESO-LoRA-R consistently outperforms these pretraining-oriented baselines in the fine-tuning regime.
F.3 Natural Language Generation: Ablation Study and Computational Cost
We conduct an ablation study on PESO-LoRA-R by fine-tuning Llama-3.1-8B and evaluating the model on GSM8K. We vary the rank $r$ and the restart frequency $K$, and additionally compare using one versus two restarts per epoch. We set the scale parameter to $\gamma=128$, and all other hyperparameters, unless otherwise noted, follow Appendix F.2. The results are presented in Table 9 and 10.
Overall, the results indicate that PESO-LoRA-R does not require a large rank to achieve strong performance: accuracy improves substantially even at small ranks, with diminishing gains beyond $r=8$. Furthermore, the method shows robust behavior across a wide range of restart frequencies, as performance remains stable when varying $K$ or the number of restarts performed per epoch.
We also report memory and runtime for Llama-3.1-8B fine-tuning on GSM8K with rank $r=8$, $K=500$, and restricting one restart per epoch. As shown in Table 11, PESO-LoRA-R matches the memory footprint of vanilla LoRA and incurs only a negligible increase in runtime. Specifically, the SVD (incurred when doing restart) step accounts for only 1.9% of the total computation time in our experiments, indicating that its overhead is negligible.
Table 9: Ablation on rank $r$ for PESO-LoRA-R, evaluated on GSM8K with restart frequency $K=500$.
Rank $r$ 1 2 4 8 16 32 64
Acc (%) 75.28 75.97 76.88 77.56 77.71 78.01 78.39
Table 10: Ablation on restart frequency $K$ and restart count (once or twice per epoch), evaluated on GSM8K with rank $r=8$. Infinity denotes applying exploration only at initialization.
$K$ 250 500 750 1000 Infinity
Restart Once 77.18 77.56 77.18 77.26 77.26
Restart Twice 76.88 77.30 77.56 76.88
Table 11: Memory (excluding transient peaks) and time for Llama-3.1-8B fine-tuning on GSM8K with $r=8$, $K=500$, and one restart per epoch.
Method LoRA PESO-LoRA-R
Time 5h03m 5h09m
Memory 25.7G 25.7G
F.4 Multi-Epoch Low-Rank Analysis: Hyperparameter Settings
In Section 4.3, we fine-tune the T5-base model on SST-2 dataset for 4 epochs. We vary the rank of LoRA in $\{2,4,8\}$, keep the rank of PESO-LoRA-R as 2, and add full-parameter fine-tuning for comparison. We keep all the other hyperparameter settings the same as in F.1.
Appendix G Asymptotic Convergence
In this section, we analyze the asymptotic convergence of the general PESO framework (Algorithm 1). To maintain generality, we do not restrict the specific choices of subspaces $\{\mathcal{S}_{k}\}$ or the subroutines Opt and UpdateM. Consequently, the derived convergence bounds may contain a bias term driven by the alignment quality between the subspace and the true gradient. We formalize this quality by assuming the subspace $\mathcal{S}_{k}$ sufficiently captures the gradient $G_{k}$ at restart indices.
Assumption 5.
There exists a sequence $\{\delta_{k}\}\geq 0$ such that $dist(G_{k},\mathcal{S}_{k})\leq\delta_{k}$ for $k$ where full gradient restart is implemented. Furthermore, $\lim_{k\to\infty}\delta_{k}<\infty$.
Remark 2.
In the general case, if $\liminf_{k\to\infty}\delta_{k}>0$, the convergence established in this section may contain a bias term depending on the approximation error. However, this bias can be eliminated in specific instantiations: (i) if $\mathcal{S}_{k}$ is chosen such that $G_{k}\in\mathcal{S}_{k}^{\text{FG}}$, then $\delta_{k}=0$ and the optimality measure converges to zero. (ii) in a deterministic setting, if $\mathcal{S}_{k}$ is chosen as the top-$r$ SVD subspace of $G_{k}$ (as in PESO-LoRA-R), the bias term vanishes. In this case, we establish exact asymptotic convergence, i.e., $\liminf_{k\to\infty}\|G_{k}\|=0$. We provide the full details in Appendix G.2. (iii) Even in the stochastic case, exact convergence can be recovered. By specifying the subspace construction as top-$r$ SVD (to control projection bias) and adopting a carefully tuned hyperparameter schedule, PESO-LoRA-R achieves exact convergence. The proof of this result (Theorem 5.1) is provided in Appendix H.
G.1 Deterministic Case
For completeness, we first consider the deterministic setting, where the gradients $G_{k}$ used by PESO (Algorithm 1) are exact and free of stochastic noise. We begin by stating a mild assumption on the behavior of Opt and UpdateM, which ensures that each iteration produces a descent step.
Assumption 6.
Opt and UpdateM generate the updates satisfying $\ell(W_{k})\leq\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))$ and $\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))\leq\ell(W_{k-1})$ for all $k=1,2,\cdots$.
Assumption 6 requires that both Opt and UpdateM are descent-preserving, ensuring that the objective value is non-increasing across iterations. This condition is standard and holds, for example, when both procedures are implemented via gradient descent with step size $\alpha\leq 1/L$, where $L$ is the Lipschitz constant of the gradient as specified in Assumption 1. In particular, UpdateM can be implemented as a warm-started gradient step on the subspace parameters with respect to the original loss function, initialized from the previous iterate. Such descent guarantees follow from classical smooth optimization results; see, e.g., Nesterov (2018).
Proposition G.1.
Let Opt and UpdateM be gradient descent schemes on $\ell$ with constant learning rate $\alpha\leq 1/L$. Then Assumption 6 holds.
We now present the deterministic convergence result.
Theorem G.2.
Suppose Assumptions 1, 5, and 6 hold. With full gradient restart enabled and learning rate $\eta_{k}=\tfrac{1}{L}$, the iterates $\{W_{k}\}$ generated by Algorithm 1 satisfy $\liminf_{k\to\infty}\|G_{k}\|\leq\lim_{k\to\infty}\delta_{k}$.
Proof.
We assume the frequency of the full gradient restart is $K$. By the descent lemma, whenever $k-1\mod K=0$ (i.e., when a full gradient restart occurs), define
$\widehat{W}_{k}:=W_{k-1}-\tfrac{1}{L}P_{\operatorname*{\mathcal{S}}_{k}}(G_{k}).$ (13)
Performing a full gradient restart as UpdateM, as discussed in (8), is thus equivalent to moving from $W_{k-1}$ to $\widehat{W}_{k}$. It follows that
$\begin{split}\ell(\widehat{W}_{k})=\ell(W_{k-1}-\tfrac{1}{L}P_{\operatorname*{\mathcal{S}}_{k}}(G_{k}))&\leq\ell(W_{k-1})+\left\langle G_{k},-\tfrac{1}{L}P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\right\rangle+\tfrac{L}{2}\|\tfrac{1}{L}P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}\\ &=\ell(W_{k-1})-\tfrac{1}{L}\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}+\tfrac{L}{2}\|\tfrac{1}{L}P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}\\ &=\ell(W_{k-1})-\tfrac{1}{2L}\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2},\end{split}$ (14)
where the second equality holds because projection onto a subspace is orthogonal.
Therefore, by Assumption 6, for iterates $i=k,\ldots,k+K-1$ (note that $\ell(W_{k})\leq\ell(\widehat{W}_{k})$ since $W_{k}$ is obtained by applying Opt to $\widehat{W}_{k}$),
$\tfrac{1}{2L}\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}\leq\ell(W_{k-1})-\ell(\widehat{W}_{k})\leq\ell(W_{k-1})-\ell(W_{i}).$ (15)
Importantly, (15) remains valid regardless of how frequently other types of UpdateM are applied between full gradient restarts, since all updates preserve the descent property by Assumption 6. In particular, whenever $\mathcal{M}_{k}$ is updated without a full gradient restart, we have
$\begin{split}\ell(W_{k-1})-\ell(W_{k})&=\ell(\widetilde{W}_{k-1}+\mathcal{M}_{k-1}(\xi_{k-1}))-\ell(W_{k})\\ &\geq\ell(\widetilde{W}_{k-1}+\mathcal{M}_{k-1}(\xi_{k-1}))-\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))\\ &\geq 0.\end{split}$ (16)
This ensures the chain of inequalities in (15) continues to hold when updates are performed by OptM.
Hence, for any integer $k\in\mathbb{N}$,
$\tfrac{1}{2L}\|P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|^{2}\leq\ell(W_{kK})-\ell(W_{(k+1)K}).$ (17)
Here $kK$ and $(k+1)K$ denote integer products.
Since $\{\ell(W_{k})\}$ is bounded below (Assumption 1) and monotonically decreasing (Assumption 6 together with the descent lemma at restart points), it converges by the monotone convergence theorem and is Cauchy. Thus $\ell(W_{k})-\ell(W_{k+1})\to 0$, and
$\tfrac{1}{2L}\|P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|^{2}\to 0.$ (18)
Finally, note that
$\|G_{kK+1}\|\leq\mathrm{dist}(G_{kK+1},\mathcal{S}_{kK+1})+\|P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|\leq\delta_{kK+1}+\|P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|\to\delta,$ (19)
where $\delta:=\lim_{k\to\infty}\delta_{k}$. Therefore, $\liminf_{k\to\infty}\|G_{k}\|\leq\delta$. ∎
G.2 Exact Convergence of Deterministic PESO-LoRA-R
In this section, we present a simple case study in which PESO-LoRA-R admits exact asymptotic convergence. We consider the deterministic setting where the subspace $\mathcal{S}_{k}^{\text{FG}}$ is chosen as the rank-$r$ subspace spanned by the top singular vectors of $G_{k}$, consistent with the construction used in PESO-LoRA-R. Under this setting, we show that the algorithm converges asymptotically to a stationary point without approximation error induced by subspace restriction. This result is consistent with the result in He et al. (2024), which establishes exact convergence for GaLore under deterministic gradients.
We first present the following lemma, which bounds the error introduced by the projected gradients.
Lemma G.3.
Let $G\in\mathbb{R}^{m\times n}$ have singular value decomposition
$G=U\Sigma V^{\top},\qquad \Sigma=\mathrm{diag}(\sigma_{1},\dots,\sigma_{p}),\quad\sigma_{1}\geq\cdots\geq\sigma_{p}\geq 0,$
where $p=\min\{m,n\}$. Let
$U_{r}:=U[:,1:r],\quad V_{r}:=V[:,1:r],\quad\Lambda_{r}:=\mathrm{diag}(\sigma_{1},\dots,\sigma_{r}),$
and define $G_{r}:=U_{r}\Lambda_{r}V_{r}^{\top}$. Then the following bounds hold:
$\|G-G_{r}\|_{F}^{2}\;\leq\;\Bigl(1-\frac{r}{p}\Bigr)\,\|G\|_{F}^{2}\quad\text{and}\quad\|G_{r}\|_{F}^{2}\;\geq\;\frac{r}{p}\,\|G\|_{F}^{2}.$ (20)
Proof.
By the Eckart–Young–Mirsky theorem,
$\|G-G_{r}\|_{F}^{2}=\sum_{i=r+1}^{p}\sigma_{i}^{2},\qquad\|G\|_{F}^{2}=\sum_{i=1}^{p}\sigma_{i}^{2},\qquad\|G_{r}\|_{F}^{2}=\sum_{i=1}^{r}\sigma_{i}^{2}.$
Since the singular values are sorted in nonincreasing order, we have
$\frac{1}{r}\sum_{i=1}^{r}\sigma_{i}^{2}\;\geq\;\frac{1}{p}\sum_{i=1}^{p}\sigma_{i}^{2},$
which implies the second inequality in (20):
$\|G_{r}\|_{F}^{2}=\sum_{i=1}^{r}\sigma_{i}^{2}\;\geq\;\frac{r}{p}\sum_{i=1}^{p}\sigma_{i}^{2}=\frac{r}{p}\,\|G\|_{F}^{2}.$
The first inequality follows immediately from the Pythagorean identity $\|G\|_{F}^{2}=\|G_{r}\|_{F}^{2}+\|G-G_{r}\|_{F}^{2}$:
$\|G-G_{r}\|_{F}^{2}=\|G\|_{F}^{2}-\|G_{r}\|_{F}^{2}\;\leq\;\|G\|_{F}^{2}-\frac{r}{p}\,\|G\|_{F}^{2}=\Bigl(1-\frac{r}{p}\Bigr)\,\|G\|_{F}^{2}.$
Next, we state the exact convergence result.
Theorem G.4.
Suppose Assumption 1 and 6 hold. Assume that full-gradient restarts are enabled, the learning rate is chosen as $\eta_{k}=\tfrac{1}{L}$, and the subspace $\mathcal{S}_{k}^{\mathrm{FG}}$ is selected as the top-$r$ SVD subspace of $G_{k}$, as in PESO-LoRA-R. Then the iterates $\{W_{k}\}$ generated by Algorithm 1 satisfy $\liminf_{k\to\infty}\|G_{k}\|=0$.
Proof.
The proof follows the same structure as that of Theorem G.2, with one key modification. Since now $\mathcal{S}_{kK+1}$ is chosen as the top-$r$ SVD subspace of $G_{kK+1}$,
$\begin{split}\|G_{kK+1}\|&=\|G_{kK+1}-P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})+P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|\\ &\leq\sqrt{1-\frac{r}{p}}\,\|G_{kK+1}\|+\|P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|,\end{split}$ (21)
where $p:=\min\{m,n\}$ and the inequality follows from Lemma G.3 and the triangle inequality. It follows from (18) that
$\left(1-\sqrt{1-\frac{r}{p}}\right)\|G_{kK+1}\|\leq\|P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|\to 0,$
and therefore $\liminf_{k\to\infty}\|G_{k}\|=0$. ∎
G.3 Stochastic Case
We now turn to the stochastic setting, where only a stochastic estimator $\widetilde{G}_{k}$ of the true gradient $G_{k}$ is available. Our goal is to establish asymptotic convergence under standard assumptions on the learning rates and the stochasticity of the updates. We begin by stating a condition on the learning rates used in the full-gradient restart step.
Assumption 7.
The learning rates for full gradient restart in (8) satisfies $\sum\eta_{k}=\infty$ and $\sum\eta_{k}^{2}<\infty$.
Assumption 7 is standard in stochastic optimization; see, e.g., Bottou et al. (2018). We next impose a stochastic descent condition on the updates generated by Opt and UpdateM.
Assumption 8.
Opt and UpdateM generate the updates satisfying $\mathbb{E}[\ell(W_{k})]\leq\mathbb{E}[\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))]+C_{k}$ and $\mathbb{E}[\ell(\widetilde{W}_{k}+\mathcal{M}_{k}(\xi_{k-1}))]\leq\mathbb{E}[\ell(W_{k-1})]+C_{k}$ where $\sum_{k}|C_{k}|<\infty$.
Assumption 8 can be viewed as a stochastic analogue of Assumption 6, allowing for controlled noise in the descent property. This assumption is mild and is satisfied, for instance, when both Opt and UpdateM are implemented using stochastic gradient descent with diminishing step sizes; see below.
We begin by verifying the validity of Assumption 8. As an illustrative case, suppose Opt and UpdateM are implemented by SGD with diminishing step sizes $\{\alpha_{k}\}$ satisfying $\sum_{k}\alpha_{k}^{2}<\infty$. Let $\widehat{W}_{k}$ denote the weight after such an update. By the descent lemma (see also (Bottou et al., 2018, Lemma 4.4)), the expected decrease can be bounded as
$\mathbb{E}[\ell(\widehat{W}_{k})]\;\leq\;\mathbb{E}[\ell(W_{k})]-\alpha_{k}\!\left(1-\tfrac{L\alpha_{k}}{2}\right)\!\mathbb{E}\|G_{k}\|^{2}+\tfrac{L}{2}\alpha_{k}^{2}C.$ (22)
For $\alpha_{k}\leq 1/L$, this simplifies to
$\mathbb{E}[\ell(\widehat{W}_{k})]-\mathbb{E}[\ell(W_{k})]\;\leq\;-\tfrac{\alpha_{k}}{2}\,\mathbb{E}\|G_{k}\|^{2}+\tfrac{L}{2}\alpha_{k}^{2}C.$ (23)
Taking positive and negative parts, we obtain
$\big[\mathbb{E}[\ell(\widehat{W}_{k})]-\mathbb{E}[\ell(W_{k})]\big]_{+}\;\leq\;\tfrac{L}{2}C\alpha_{k}^{2},\qquad\big[\mathbb{E}[\ell(W_{k})]-\mathbb{E}[\ell(\widehat{W}_{k})]\big]_{+}\;\leq\;\mathbb{E}[\ell(W_{k})]-\mathbb{E}[\ell(\widehat{W}_{k})].$ (24)
Summing over all $k$,
$\begin{split}\sum_{k=0}^{\infty}\big[\mathbb{E}[\ell(\widehat{W}_{k})]-\mathbb{E}[\ell(W_{k})]\big]_{+}&\;\leq\;\tfrac{L}{2}C\sum_{k=0}^{\infty}\alpha_{k}^{2}\;<\;\infty,\\ \sum_{k=0}^{\infty}\big[\mathbb{E}[\ell(W_{k})]-\mathbb{E}[\ell(\widehat{W}_{k})]\big]_{+}&\;\leq\;\sup_{k}\mathbb{E}[\ell(W_{k})]-\ell(W^{*})\;<\;\infty.\end{split}$ (25)
Defining $C_{k}:=\mathbb{E}[\ell(\widehat{W}_{k})]-\mathbb{E}[\ell(W_{k})]$, we conclude that $\sum_{k}|C_{k}|<\infty$, hence Assumption 8 holds.
We are now ready to present the proof of our main result, Theorem G.5. Importantly, the result holds for any choice of UpdateM, whether warm-start or restart, and at any frequency, as long as the assumptions are satisfied. In particular, Algorithm 1 may combine different UpdateM strategies at varying frequencies, and the guarantee remains valid.
Theorem G.5.
Suppose Assumption 12578 hold. In addition, assume
$\mathbb{E}[\widetilde{G}_{k}\mid W_{k-1}]=G_{k}\qquad\text{for all }k,$ (26)
and that $\operatorname*{\mathcal{S}}_{k}$ is measurable with respect to $W_{k-1}$. With full gradient restart, the iterates $\{W_{k}\}$ generated by Algorithm 1 satisfy
$\liminf_{k\to\infty}\mathbb{E}[\|G_{k}\|]\leq\lim_{k\to\infty}\delta_{k}.$
Proof.
Because $\operatorname*{\mathcal{S}}_{k}$ are subspaces and by the measurability assumption, $P_{\operatorname*{\mathcal{S}}_{k}}$ is fixed given $W_{k-1}$. Therefore, one has
$\begin{split}\mathbb{E}\big(\|P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\|^{2}\big)&=\mathbb{E}\Big[\mathbb{E}\big(\|P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\|^{2}\mid W_{k-1}\big)\Big]\\ &=\mathbb{E}\Big[\big\|\mathbb{E}\big(P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\mid W_{k-1}\big)\big\|^{2}+\mathbb{V}\big(P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\mid W_{k-1}\big)\Big]\\ &=\mathbb{E}\Big[\|P_{\operatorname*{\mathcal{S}}_{k}}(\mathbb{E}[\widetilde{G}_{k}\mid W_{k-1}])\|^{2}+\mathbb{E}\big(\|P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k}-\mathbb{E}[\widetilde{G}_{k}\mid W_{k-1}])\|^{2}\mid W_{k-1}\big)\Big]\\ &=\mathbb{E}\Big[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}+\mathbb{E}\big(\|P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k}-G_{k})\|^{2}\mid W_{k-1}\big)\Big]\\ &\leq\mathbb{E}\big[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}\big]+\mathbb{E}\big[\|\widetilde{G}_{k}-G_{k}\|^{2}\big]\\ &\leq\mathbb{E}\big[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}\big]+C.\end{split}$ (27)
Here, $C$ comes from the bound in Assumption 2.
Similar to the deterministic case, we assume the frequency of the full gradient restart is $K$, and consider $k-1\mod K=0$ (i.e., when a full gradient restart occurs). Similarly, we define
$\widehat{W}_{k}:=W_{k-1}-\eta_{k}P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k}).$ (28)
By the property of the full gradient restart, we have
$\begin{split}\mathbb{E}[\ell(\widehat{W}_{k})]&\leq\mathbb{E}[\ell(W_{k-1})]+\mathbb{E}\big[\langle G_{k},-\eta_{k}P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\rangle\big]+\frac{L}{2}\mathbb{E}[\|\eta_{k}P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\|^{2}]\\ &=\mathbb{E}[\ell(W_{k-1})]-\eta_{k}\,\mathbb{E}\Big[\big\langle P_{\operatorname*{\mathcal{S}}_{k}}(G_{k}),\,\mathbb{E}[P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\mid W_{k-1}]\big\rangle\Big]+\frac{L\eta_{k}^{2}}{2}\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\|^{2}]\\ &=\mathbb{E}[\ell(W_{k-1})]-\eta_{k}\,\mathbb{E}\big[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}\big]+\frac{L\eta_{k}^{2}}{2}\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\|^{2}]\\ &\leq\mathbb{E}[\ell(W_{k-1})]-\eta_{k}\,\mathbb{E}\big[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}\big]+\frac{L\eta_{k}^{2}}{2}\Big(\mathbb{E}\big[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}\big]+C\Big)\\ &=\mathbb{E}[\ell(W_{k-1})]-\left(\eta_{k}-\frac{L\eta_{k}^{2}}{2}\right)\mathbb{E}\big[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}\big]+\frac{L\eta_{k}^{2}}{2}C,\end{split}$ (29)
where the first inequality follows from Assumption 1, the second equality uses linearity of projection and the tower property, and the third equality uses
$\mathbb{E}[P_{\operatorname*{\mathcal{S}}_{k}}(\widetilde{G}_{k})\mid W_{k-1}]=P_{\operatorname*{\mathcal{S}}_{k}}(\mathbb{E}[\widetilde{G}_{k}\mid W_{k-1}])=P_{\operatorname*{\mathcal{S}}_{k}}(G_{k}),$
which follows from (26) and measurability of $P_{\operatorname*{\mathcal{S}}_{k}}$.
Then by Assumption 8, suppose $k-1\mod{K}=0$, and for iterates $i=k+1,\cdots,k+K-1$,
$\mathbb{E}[\ell(W_{i})]\leq\mathbb{E}[\ell(W_{i-1})]+2C_{i},$ (30)
where $2C_{i}$ comes from bounding the scenario where both Opt and UpdateM operate at $i$-th iterate. Then summing up the inequalities (29) and (30) for $i=k+1,\cdots,k+K-1$, and use the fact that $\mathbb{E}[\ell(W_{k})]\leq\mathbb{E}[\ell(\widehat{W}_{k})]+C_{k}$ since $W_{k}$ is obtained by applying Opt to $\widehat{W}_{k}$, we have
$\left(\eta_{k}-\frac{L\eta_{k}^{2}}{2}\right)\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}]-\frac{L\eta_{k}^{2}}{2}C-C_{k}-2\sum_{i=k+1}^{k+K-1}C_{i}\leq\mathbb{E}[\ell(W_{k-1})-\ell(W_{k+K-1})].$ (31)
By Assumption 7, $\eta_{k}\to 0$ so without loss of generality, we can assume $\frac{L\eta_{k}}{2}\leq\frac{1}{2}$ for any $k\in\mathbb{N}$. Therefore for all integer $k\in\mathbb{N}$, we have
$\frac{\eta_{kK+1}}{2}\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|^{2}]-\frac{L\eta_{kK+1}^{2}}{2}C-C_{kK+1}-2\sum_{i=kK+2}^{(k+1)K}C_{i}\leq\mathbb{E}[\ell(W_{kK})-\ell(W_{(k+1)K})].$ (32)
By Assumption 1, there exists a constant $\ell_{\star}$ so that $\ell_{\star}\leq\ell(W)$ for any $W$. Summing up for $k\in\{1,\cdots,T\}$ one has
$\begin{split}\sum_{k=1}^{T}\frac{\eta_{kK+1}}{2}\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|^{2}]\leq&\frac{LC}{2}\sum_{k=1}^{T}\eta_{kK+1}^{2}+\sum_{k=1}^{T}C_{kK+1}+2\sum_{k=1}^{T}\sum_{i=kK+2}^{(k+1)K}C_{i}\\ &+\mathbb{E}[\ell(W_{K})]-\ell_{\star}.\end{split}$ (33)
Taking $T\to\infty$, and note that $\sum_{k}\eta_{k}^{2}<\infty$ and $\sum_{k}|C_{k}|<\infty$, the first and second series on the right hand side of (33) are obviously bounded. For the third series, note that
$\sum_{k=1}^{\infty}\sum_{i=kK+2}^{(k+1)K}C_{i}\leq\sum_{k=1}^{\infty}\sum_{i=kK+2}^{(k+1)K}|C_{i}|\leq\sum_{i=1}^{\infty}|C_{i}|<\infty.$ (34)
Finally $\|\mathbb{E}[\ell(W_{K})]$-$\ell_{\star}\|$ is obviously bounded for a fixed $K$, and therefore, one has
$\sum_{k=1}^{\infty}\eta_{kK+1}\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{kK+1}}(G_{kK+1})\|^{2}]<\infty.$ (35)
By $\sum_{k}^{\infty}\eta_{k}=\infty$ and a contradiction argument, one has
$\liminf_{k\to\infty}\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|^{2}]=0\quad\Rightarrow\quad\liminf_{k\to\infty}\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|]=0,$
where the implication follows from Jensen’s inequality $\mathbb{E}\|X\|\leq\sqrt{\mathbb{E}\|X\|^{2}}$.
Since
$\|G_{k}\|\leq\operatorname{dist}(G_{k},\mathcal{S}_{k})+\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|\leq\delta_{k}+\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|,$
taking expectations and $\liminf$ yields
$\liminf_{k\to\infty}\mathbb{E}[\|G_{k}\|]\leq\lim_{k\to\infty}\delta_{k}+\liminf_{k\to\infty}\mathbb{E}[\|P_{\operatorname*{\mathcal{S}}_{k}}(G_{k})\|]=\lim_{k\to\infty}\delta_{k},$
which completes the proof. ∎
Appendix H Non-asymptotic Rate: Proof of Theorem 5.1
Algorithm 7 PESO-LoRA-R (Analysis Version)
1:Pre-trained parameters $W_{0}\in\mathbb{R}^{m\times n}$, frequency $K$, rank $r$.
2:Restart stepsizes $\{\eta_{k}\}$ (used only at restart steps), Adam stepsizes $\{\alpha_{k}\}$.
3:Adam hyperparameters $(\beta_{1},\beta_{2},\epsilon)$.
4:Set $k\leftarrow 1$, $\widetilde{W}_{0}\leftarrow W_{0}$, $A_{0}\leftarrow 0$, $B_{0}\leftarrow 0$.
5:Initialize Adam states $m_{A}\leftarrow 0$, $v_{A}\leftarrow 0$, $m_{B}\leftarrow 0$, $v_{B}\leftarrow 0$.
6:while stopping criteria not satisfied do
7:  if $(k-1)\bmod K=0$ then
8:   $\widetilde{W}_{k}\leftarrow\widetilde{W}_{k-1}+A_{k-1}B_{k-1}$.
9:   Compute stochastic full gradient estimator $\widetilde{G}_{k}$ at $\widetilde{W}_{k}$.
10:   $(U_{k},\Lambda_{k},V_{k})\leftarrow{\tt SVD}(-\widetilde{G}_{k})$
11:   Set $A_{k-1}\leftarrow\sqrt{\eta_{k}}\,U_{k}\sqrt{\Lambda_{k}}$,  $B_{k-1}\leftarrow\sqrt{\eta_{k}}\,\sqrt{\Lambda_{k}}V_{k}^{\top}$.
12:   Reset Adam states: $m_{A},v_{A},m_{B},v_{B}\leftarrow 0$.
13:  else
14:   $\widetilde{W}_{k}\leftarrow\widetilde{W}_{k-1}$.
15:  end if
16:  Compute stochastic gradients $g_{A,k}\leftarrow\nabla_{A}\ell(\widetilde{W}_{k}+A_{k-1}B_{k-1})$, $g_{B,k}\leftarrow\nabla_{B}\ell(\widetilde{W}_{k}+A_{k-1}B_{k-1})$.
17:  $m_{A}\leftarrow\beta_{1}m_{A}+(1-\beta_{1})g_{A,k}$, $v_{A}\leftarrow\beta_{2}v_{A}+(1-\beta_{2})g_{A,k}\odot g_{A,k}$.
18:  $m_{B}\leftarrow\beta_{1}m_{B}+(1-\beta_{1})g_{B,k}$, $v_{B}\leftarrow\beta_{2}v_{B}+(1-\beta_{2})g_{B,k}\odot g_{B,k}$.
19:  Bias corrections: $\widehat{m}_{A}\leftarrow m_{A}/(1-\beta_{1}^{k})$, $\widehat{v}_{A}\leftarrow v_{A}/(1-\beta_{2}^{k})$, and similarly for $(\widehat{m}_{B},\widehat{v}_{B})$.
20:  Adam updates: $A_{k}\leftarrow A_{k-1}-\alpha_{k}\,\widehat{m}_{A}/(\sqrt{\widehat{v}_{A}}+\epsilon)$.
21:  $B_{k}\leftarrow B_{k-1}-\alpha_{k}\,\widehat{m}_{B}/(\sqrt{\widehat{v}_{B}}+\epsilon)$.
22:  $k\leftarrow k+1$.
23:end while
24:return $\widetilde{W}_{k-1}+A_{k-1}B_{k-1}$.
In this section, we provide the proof of the non-asymptotic convergence rate stated in Theorem 5.1 for PESO-LoRA-R. To facilitate the analysis, we focus on Algorithm 7, which simplifies the practical implementation (Algorithm 5). Specifically, Algorithm 7 substitutes AdamW with standard Adam and employs an explicit reset of the momentum and velocity states at each restart, replacing the smoothing and alignment heuristics described in Appendix D.
We follow the notation in the main text: $k$ denotes the global iteration index, and a restart happens when $(k-1)\bmod K=0$. For conciseness in the proof, we index restart cycles by $s\in\{0,1,\dots\}$ and denote the restart iteration by
$k_{s}:=sK+1.$
We write $\widetilde{W}_{s}:=W_{k_{s}}$ for the weights before doing the restart $s$ (which is the updated anchored state). Within cycle $s$, PESO-LoRA-R performs $K$ inner Adam steps producing $\widetilde{W}_{s+1}$ at the next restart. We assume Assumptions 14 from the main text. For clarity, we restate Assumption 3 and 4 that require careful measurability.
Assumption 9 (Detailed version of Assumption 4).
Let $P_{r}(\cdot)$ denote the rank-$r$ truncated SVD operator mapping a matrix to its best rank-$r$ approximation in Frobenius norm. Assume there exists a sequence $\{\varepsilon_{s}\}_{s\geq 0}$ such that at each restart iterate $k_{s}=sK+1$,
$\big\|\mathbb{E}\!\left[P_{r}(\widetilde{G}_{k_{s}})\mid\widetilde{W}_{s}\right]-P_{r}(G_{k_{s}})\big\|_{F}\leq\varepsilon_{s},\quad\text{and}\quad\sum_{s=0}^{\infty}\eta_{s}\,\varepsilon_{s}^{2}<\infty,$ (36)
where $\eta_{s}$ is the restart stepsize at cycle $s$.
Assumption 10 (Detailed version of Assumption 3).
For each restart cycle $s$, define the inner objective $f_{s}(A,B):=\ell(\widetilde{W}_{s}+AB)$ and let $X:=(A,B)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{r\times n}$ with dimension $d:=r(m+n)$. Let $\{X_{s,t}\}_{t=0}^{K}$ be generated by Adam on $f_{s}$ for $K$ steps with per-cycle reset $m_{s,0}=0$, $v_{s,0}=0$. At each inner step $t\geq 1$, let $g_{s,t}$ be the stochastic gradient (with respect to $X$) used by Adam. Let $\mathcal{F}_{s,t-1}$ denote the filtration generated by all random variables (gradients and iterates) up to inner step t1t-1 of cycle $s$. Assume:
  1. (i)
    $f_{s}$ has $L_{\rm in}$-Lipschitz continuous gradients;
  2. (ii)
    $\mathbb{E}[g_{s,t}\mid\mathcal{F}_{s,t-1}]=\nabla f_{s}(X_{s,t-1})$;
  3. (iii)
    $\|g_{s,t}\|_{\infty}\leq G_{\infty}$ almost surely.
The proof proceeds in three parts: first, we establish the descent achieved by the restart step (Lemma H.1); next, we bound the progress of the inner Adam updates on $X$ (Lemma H.2); and finally, we combine these results to derive the convergence rate (Theorem H.3).
We begin with deriving a descent lemma for the outer restart update. At restart $k_{s}=sK+1$, define the post-restart point
$\widehat{W}_{s}\;:=\;\widetilde{W}_{s}-\eta_{s}\,P_{r}(\widetilde{G}_{k_{s}}).$ (37)
Lemma H.1 (Descent at restart).
Let $\rho:=r/p$ from Lemma G.3. Under Assumptions 1, 2, and 9, for any $s\geq 0$,
$\mathbb{E}\big[\ell(\widehat{W}_{s})\big]\;\leq\;\mathbb{E}\big[\ell(\widetilde{W}_{s})\big]-\Bigl(\rho\eta_{s}-{L\eta_{s}^{2}}\Bigr)\,\mathbb{E}\|G_{k_{s}}\|_{F}^{2}+\eta_{s}\,\mathbb{E}\big[\|G_{k_{s}}\|_{F}\,\varepsilon_{s}\big]+{L\eta_{s}^{2}}\,C.$ (38)
Proof.
Fix $s$ and condition on $\widetilde{W}_{s}$. By $L$-smoothness of $\ell$ and the update rule (37),
$\ell(\widehat{W}_{s})\leq\ell(\widetilde{W}_{s})+\left\langle\nabla\ell(\widetilde{W}_{s}),\,\widehat{W}_{s}-\widetilde{W}_{s}\right\rangle+\frac{L}{2}\|\widehat{W}_{s}-\widetilde{W}_{s}\|_{F}^{2}=\ell(\widetilde{W}_{s})-\eta_{s}\langle G_{k_{s}},\,P_{r}(\widetilde{G}_{k_{s}})\rangle+\frac{L\eta_{s}^{2}}{2}\|P_{r}(\widetilde{G}_{k_{s}})\|_{F}^{2}.$
Taking the conditional expectation given $\widetilde{W}_{s}$ yields
$\mathbb{E}\big[\ell(\widehat{W}_{s})\mid\widetilde{W}_{s}\big]\leq\ell(\widetilde{W}_{s})-\eta_{s}\left\langle G_{k_{s}},\,\mathbb{E}[P_{r}(\widetilde{G}_{k_{s}})\mid\widetilde{W}_{s}]\right\rangle+\frac{L\eta_{s}^{2}}{2}\,\mathbb{E}\big[\|P_{r}(\widetilde{G}_{k_{s}})\|_{F}^{2}\mid\widetilde{W}_{s}\big].$ (39)
We first lower-bound the descent term:
$\begin{split}\left\langle G_{k_{s}},\,\mathbb{E}[P_{r}(\widetilde{G}_{k_{s}})\mid\widetilde{W}_{s}]\right\rangle&=\langle G_{k_{s}},\,P_{r}(G_{k_{s}})\rangle+\left\langle G_{k_{s}},\,\mathbb{E}[P_{r}(\widetilde{G}_{k_{s}})\mid\widetilde{W}_{s}]-P_{r}(G_{k_{s}})\right\rangle\\ &\geq\|P_{r}(G_{k_{s}})\|_{F}^{2}-\|G_{k_{s}}\|_{F}\big\|\mathbb{E}[P_{r}(\widetilde{G}_{k_{s}})\mid\widetilde{W}_{s}]-P_{r}(G_{k_{s}})\big\|_{F}\\ &\geq\|P_{r}(G_{k_{s}})\|_{F}^{2}-\|G_{k_{s}}\|_{F}\,\varepsilon_{s}.\end{split}$ (40)
where the first inequality uses $\langle G,P_{r}(G)\rangle=\|P_{r}(G)\|^{2}$ and Cauchy–Schwarz, and the second follows from Assumption 9. Multiplying by $\eta_{s}$ reverses the inequality, yielding
$\eta_{s}\left\langle G_{k_{s}},\,\mathbb{E}[P_{r}(\widetilde{G}_{k_{s}})\mid\widetilde{W}_{s}]\right\rangle\leq-\eta_{s}\|P_{r}(G_{k_{s}})\|_{F}^{2}+\eta_{s}\|G_{k_{s}}\|_{F}\,\varepsilon_{s}.$ (41)
Next, we bound the quadratic term:
$\begin{split}\frac{L\eta_{s}^{2}}{2}\,\mathbb{E}\big[\|P_{r}(\widetilde{G}_{k_{s}})\|_{F}^{2}\mid\widetilde{W}_{s}\big]&\leq\frac{L\eta_{s}^{2}}{2}\,\mathbb{E}\big[\|\widetilde{G}_{k_{s}}\|_{F}^{2}\mid\widetilde{W}_{s}\big]\\ &\leq\frac{L\eta_{s}^{2}}{2}\,\left(2\|G_{k_{s}}\|_{F}^{2}+2\mathbb{E}\big[\|\widetilde{G}_{k_{s}}-G_{k_{s}}\|_{F}^{2}\mid\widetilde{W}_{s}\big]\right)\\ &\leq\frac{L\eta_{s}^{2}}{2}\,\left(2\|G_{k_{s}}\|_{F}^{2}+2C\right)\end{split}$ (42)
where the first inequality is from $\|P_{r}(X)\|_{F}\leq\|X\|_{F}$, the second is from Young’s inequality, and the third is from Assumption 2. Plugging (41) and (42) into (39), we obtain
$\begin{split}\mathbb{E}\big[\ell(\widehat{W}_{s})\mid\widetilde{W}_{s}\big]&\leq\ell(\widetilde{W}_{s})-\eta_{s}\|P_{r}(G_{k_{s}})\|_{F}^{2}+\eta_{s}\|G_{k_{s}}\|_{F}\,\varepsilon_{s}+\frac{L\eta_{s}^{2}}{2}\left(2\|G_{k_{s}}\|_{F}^{2}+2C\right)\\ &\leq\ell(\widetilde{W}_{s})-\Bigl(\rho\eta_{s}-{L\eta_{s}^{2}}\Bigr)\|G_{k_{s}}\|_{F}^{2}+\eta_{s}\|G_{k_{s}}\|_{F}\,\varepsilon_{s}+{L\eta_{s}^{2}}C,\end{split}$ (43)
where the second inequality follows from Lemma G.3. Taking the total expectation completes the proof. ∎
We next show the progress made by Adam within one cycle. Fix a cycle $s$, the inner Adam procedure starts from $X_{s,0}=(A_{s,0},B_{s,0})$ and after $K$ Adam steps yields $X_{s,K}$. Following the indexing notation using $s$, the Adam updates take the form of
$m_{s,t}=\beta_{1,t}m_{s,t-1}+(1-\beta_{1,t})g_{s,t},\qquad v_{s,t}=\beta_{2}v_{s,t-1}+(1-\beta_{2})g_{s,t}^{\odot 2},$
with bias corrections $\widehat{m}_{s,t}=m_{s,t}/(1-\prod_{j=1}^{t}\beta_{1,j})$, $\widehat{v}_{s,t}=v_{s,t}/(1-\beta_{2}^{t})$, and
$X_{s,t}=X_{s,t-1}-\alpha_{s,t}\frac{\widehat{m}_{s,t}}{\sqrt{\widehat{v}_{s,t}}+\epsilon},\qquad t=1,\dots,K.$
Lemma H.2 (Adam progress bound).
Assume Assumption 10. Fix a restart cycle $s$ and run $K$ steps of Adam with stepsizes $\alpha_{s,t}=a_{s}/\sqrt{t}$ for $t=1,\dots,K$, initialization $m_{s,0}=v_{s,0}=0$, and hyperparameters $\beta_{2}\in(0,1)$, $\beta_{1,t}\in[0,\bar{\beta}_{1}]$ with $\bar{\beta}_{1}\in[0,1)$. Then
$\mathbb{E}\big[f_{s}(X_{s,K})\big]\leq\mathbb{E}\big[f_{s}(X_{s,0})\big]+A_{1}\,a_{s}+A_{2}\,a_{s}^{2},$ (44)
where the constants are defined by
$A_{1}:=\frac{2d\,G_{\infty}^{2}}{\epsilon}\sum_{t=1}^{K}\frac{1}{\sqrt{t}},\qquad A_{2}:=\frac{L_{\rm in}}{2}B_{1}\sum_{t=1}^{K}\frac{1}{t}$ (45)
with $B_{1}:=\frac{d}{\epsilon^{2}}G_{\infty}^{2}$.
Proof.
Let the preconditioned step be defined as $D_{s,t}:=\widehat{m}_{s,t}/(\sqrt{\widehat{v}_{s,t}}+\epsilon)$ so that $X_{s,t}=X_{s,t-1}-\alpha_{s,t}D_{s,t}$. By $L_{\rm in}$-smoothness of $f_{s}$ and Taylor expansion
$f_{s}(X_{s,t})\leq f_{s}(X_{s,t-1})-\alpha_{s,t}\langle\nabla f_{s}(X_{s,t-1}),\,D_{s,t}\rangle+\frac{L_{\rm in}}{2}\alpha_{s,t}^{2}\|D_{s,t}\|_{2}^{2}.$ (46)
Taking total expectation and summing over $t=1,\dots,K$, we have
$\mathbb{E}[f_{s}(X_{s,K})]\leq\mathbb{E}[f_{s}(X_{s,0})]-\sum_{t=1}^{K}\alpha_{s,t}\,\mathbb{E}\langle\nabla f_{s}(X_{s,t-1}),D_{s,t}\rangle+\frac{L_{\rm in}}{2}\sum_{t=1}^{K}\alpha_{s,t}^{2}\,\mathbb{E}\|D_{s,t}\|_{2}^{2}.$ (47)
We first derive a bound for $\mathbb{E}\|D_{s,t}\|_{2}^{2}$. We note that the bias-corrected momentum is $\widehat{m}_{s,t}=\sum_{\tau=1}^{t}w_{t,\tau}g_{s,\tau}$ with $\sum_{\tau=1}^{t}w_{t,\tau}=1$ and $w_{t,\tau}\geq 0$, where
$w_{t,\tau}:=\frac{(1-\beta_{1,\tau})\prod_{j=\tau+1}^{t}\beta_{1,j}}{1-\prod_{j=1}^{t}\beta_{1,j}}\geq 0.$
Therefore $\widehat{m}_{s,t}$ is a convex combination of past gradients and by Jensen’s inequality and Assumption 10(iii)
$\mathbb{E}\|\widehat{m}_{s,t}\|_{\infty}^{2}\leq\sum_{\tau=1}^{t}w_{t,\tau}\mathbb{E}\|g_{s,\tau}\|_{\infty}^{2}\leq G_{\infty}^{2}.$
Since $\sqrt{\widehat{v}_{s,t}}+\epsilon\geq\epsilon$ elementwise we then have
$\mathbb{E}\|D_{s,t}\|_{2}^{2}=\mathbb{E}\sum_{i=1}^{d}\frac{\widehat{m}_{s,t,i}^{2}}{(\sqrt{\widehat{v}_{s,t,i}}+\epsilon)^{2}}\leq\frac{1}{\epsilon^{2}}\mathbb{E}\|\widehat{m}_{s,t}\|_{2}^{2}\leq\frac{d}{\epsilon^{2}}\mathbb{E}\|\widehat{m}_{s,t}\|_{\infty}^{2}\leq B_{1}.$ (48)
We then denote the preconditioner as $H_{s,t}:=\mathrm{diag}(\sqrt{\widehat{v}_{s,t}}+\epsilon)^{-1}$. Consider the inner product term $\langle\nabla f_{s}(X_{s,t-1}),D_{s,t}\rangle$ and note that
$\langle\nabla f_{s}(X_{s,t-1}),D_{s,t}\rangle=\underbrace{\langle\nabla f_{s}(X_{s,t-1}),H_{s,t}\nabla f_{s}(X_{s,t-1})\rangle}_{(I)}+\underbrace{\langle\nabla f_{s}(X_{s,t-1}),H_{s,t}(\widehat{m}_{s,t}-\nabla f_{s}(X_{s,t-1}))\rangle}_{(II)}.$ (49)
Since $H_{s,t}\succ 0$, $(I)\geq 0$, implying $\alpha_{s,t}\mathbb{E}[(I)]\leq 0$.
We next show $\|\nabla f_{s}\|_{2}\leq\sqrt{d}G_{\infty}$. Let $\mathcal{F}_{s,t-1}$ denote the filtration generated by all random variables (gradients and iterates) up to inner step t1t-1 of cycle $s$. By Assumption 10(ii), we have $\nabla f_{s}(X_{s,t-1})=\mathbb{E}[g_{s,t}\mid\mathcal{F}_{s,t-1}]$. Using Jensen’s inequality for the convex norm $\|\cdot\|_{\infty}$ and the almost sure bound $\|g_{s,t}\|_{\infty}\leq G_{\infty}$ from Assumption 10(iii),
$\|\nabla f_{s}(X_{s,t-1})\|_{\infty}=\|\mathbb{E}[g_{s,t}\mid\mathcal{F}_{s,t-1}]\|_{\infty}\leq\mathbb{E}[\|g_{s,t}\|_{\infty}\mid\mathcal{F}_{s,t-1}]\leq G_{\infty}.$
Since $\|v\|_{2}\leq\sqrt{d}\|v\|_{\infty}$ for any $v\in\mathbb{R}^{d}$, we have
$\|\nabla f_{s}(X_{s,t-1})\|_{2}\leq\sqrt{d}\,\|\nabla f_{s}(X_{s,t-1})\|_{\infty}\leq\sqrt{d}G_{\infty}.$ (50)
For $(II)$, using Cauchy–Schwarz, $\|H_{s,t}\|_{2}\leq 1/\epsilon$, and $\|\nabla f_{s}\|_{2}\leq\sqrt{d}G_{\infty}$,
$|(II)|\leq\|\nabla f_{s}(X_{s,t-1})\|_{2}\|H_{s,t}\|_{2}\|\widehat{m}_{s,t}-\nabla f_{s}(X_{s,t-1})\|_{2}\leq\frac{\sqrt{d}G_{\infty}}{\epsilon}\|\widehat{m}_{s,t}-\nabla f_{s}(X_{s,t-1})\|_{2}.$ (51)
Moreover, by $\|\widehat{m}_{s,t}\|_{2}\leq\sqrt{d}\|\widehat{m}_{s,t}\|_{\infty}\leq\sqrt{d}G_{\infty}$ and $\|\nabla f_{s}(X_{s,t-1})\|_{2}\leq\sqrt{d}G_{\infty}$, we have
$\|\widehat{m}_{s,t}-\nabla f_{s}(X_{s,t-1})\|_{2}\leq\|\widehat{m}_{s,t}\|_{2}+\|\nabla f_{s}(X_{s,t-1})\|_{2}\leq 2\sqrt{d}\,G_{\infty}.$
Thus, we obtain a bound linear in $\alpha_{s,t}$:
$\alpha_{s,t}\mathbb{E}[(II)]\leq\alpha_{s,t}\,\frac{\sqrt{d}G_{\infty}}{\epsilon}\,(2\sqrt{d}\,G_{\infty})=\alpha_{s,t}\,\frac{2d\,G_{\infty}^{2}}{\epsilon}.$ (52)
Finally, substituting $\alpha_{s,t}=a_{s}/\sqrt{t}$ into (47), we have
$\begin{split}\mathbb{E}[f_{s}(X_{s,K})]&\leq\mathbb{E}[f_{s}(X_{s,0})]+\sum_{t=1}^{K}\alpha_{s,t}\,\frac{2d\,G_{\infty}^{2}}{\epsilon}+\frac{L_{\rm in}}{2}\sum_{t=1}^{K}\alpha_{s,t}^{2}B_{1}\\ &=\mathbb{E}[f_{s}(X_{s,0})]+a_{s}\left(\frac{2d\,G_{\infty}^{2}}{\epsilon}\sum_{t=1}^{K}\frac{1}{\sqrt{t}}\right)+a_{s}^{2}\left(\frac{L_{\rm in}}{2}B_{1}\sum_{t=1}^{K}\frac{1}{t}\right)\\ &=\mathbb{E}[f_{s}(X_{s,0})]+A_{1}a_{s}+A_{2}a_{s}^{2},\end{split}$ (53)
according to the definition of $A_{1}$ and $A_{2}$. This finishes the proof. ∎
Theorem H.3 (Convergence rate in restart cycles).
Let $\ell_{\star}:=\inf_{W}\ell(W)>$-$\infty$ and $\rho=r/p$ from Lemma G.3. Suppose Assumptions 1, 2, 9, and 10 hold. Fix the total number of restart cycles $S\geq 1$, and choose the step sizes as $\eta_{s}=\frac{1}{\sqrt{S}}$ and $a_{s}=\frac{1}{S}$. Set the hyperparameters of Adam as those in Lemma H.2. Then, the minimum expected gradient norm over the restart points satisfies
$\min_{0\leq s\leq S-1}\mathbb{E}\|G_{k_{s}}\|_{F}^{2}\;\leq\;\frac{1}{\sqrt{S}}\left(\frac{4}{\rho}\left(\mathbb{E}[\ell(\widetilde{W}_{0})]-\ell_{\star}+\frac{1}{\rho}\bar{E}_{\rm sq}+LC+A_{1}+A_{2}\right)\right),$ (54)
where $\bar{E}_{\rm sq}:=\sum_{s=0}^{S-1}\eta_{s}\varepsilon_{s}^{2}$ (finite because of Assumption 9), $C$ from Assumption 2, and $A_{1},A_{2}$ from Lemma H.2.
Proof.
We combine the descent progress from the restart step (Lemma H.1) with the drift bound from the Adam inner steps (Lemma H.2). By Lemma H.1, the update at restart $s$ satisfies:
$\mathbb{E}[\ell(\widehat{W}_{s})]\leq\mathbb{E}[\ell(\widetilde{W}_{s})]-\left(\rho\eta_{s}-L\eta_{s}^{2}\right)\mathbb{E}\|G_{k_{s}}\|_{F}^{2}+\eta_{s}\,\mathbb{E}[\|G_{k_{s}}\|_{F}\varepsilon_{s}]+L\eta_{s}^{2}C.$ (55)
Without loss of generality, we assume $\eta_{s}\leq\rho/2L$ (which holds for large $S$). Therefore,
$\rho\eta_{s}-L\eta_{s}^{2}\;\geq\;\frac{\rho}{2}\eta_{s}.$
Substituting this into (55) gives
$\mathbb{E}[\ell(\widehat{W}_{s})]\leq\mathbb{E}[\ell(\widetilde{W}_{s})]-\frac{\rho}{2}\eta_{s}\,\mathbb{E}\|G_{k_{s}}\|_{F}^{2}+\eta_{s}\,\mathbb{E}[\|G_{k_{s}}\|_{F}\varepsilon_{s}]+L\eta_{s}^{2}C.$ (56)
By Young’s inequality ($xy\leq\frac{\rho}{4}x^{2}+\frac{1}{\rho}y^{2}$ with $x=\sqrt{\eta_{s}}\|G_{k_{s}}\|_{F}$ and $y=\sqrt{\eta_{s}}\varepsilon_{s}$),
$\mathbb{E}[\ell(\widehat{W}_{s})]\leq\mathbb{E}[\ell(\widetilde{W}_{s})]-\frac{\rho}{4}\eta_{s}\,\mathbb{E}\|G_{k_{s}}\|_{F}^{2}+\frac{1}{\rho}\eta_{s}\varepsilon_{s}^{2}+L\eta_{s}^{2}C.$ (57)
By Lemma H.2, identifying $f_{s}(X_{s,0})\equiv\ell(\widehat{W}_{s})$ and $f_{s}(X_{s,K})\equiv\ell(\widetilde{W}_{s+1})$,
$\mathbb{E}[\ell(\widetilde{W}_{s+1})]\leq\mathbb{E}[\ell(\widehat{W}_{s})]+A_{1}a_{s}+A_{2}a_{s}^{2}.$ (58)
Combining these yields the recurrence for cycle $s$:
$\mathbb{E}[\ell(\widetilde{W}_{s+1})]\leq\mathbb{E}[\ell(\widetilde{W}_{s})]-\frac{\rho}{4}\eta_{s}\,\mathbb{E}\|G_{k_{s}}\|_{F}^{2}+\frac{1}{\rho}\eta_{s}\varepsilon_{s}^{2}+L\eta_{s}^{2}C+A_{1}a_{s}+A_{2}a_{s}^{2}.$ (59)
Rearranging and summing over $s=0,\dots,S-1$ we have
$\frac{\rho}{4}\sum_{s=0}^{S-1}\eta_{s}\mathbb{E}\|G_{k_{s}}\|_{F}^{2}\leq\mathbb{E}[\ell(\widetilde{W}_{0})]-\mathbb{E}[\ell(\widetilde{W}_{S})]+\frac{1}{\rho}\sum_{s=0}^{S-1}\eta_{s}\varepsilon_{s}^{2}+LC\sum_{s=0}^{S-1}\eta_{s}^{2}+\sum_{s=0}^{S-1}(A_{1}a_{s}+A_{2}a_{s}^{2}).$ (60)
We now substitute the specific step sizes $\eta_{s}=\frac{1}{\sqrt{S}}$ and $a_{s}=\frac{1}{S}$. For the Adam error terms we have
$\sum_{s=0}^{S-1}(A_{1}a_{s}+A_{2}a_{s}^{2})=A_{1}+\frac{A_{2}}{S}\leq A_{1}+A_{2}.$
Substituting this back, and by $\ell(\widetilde{W}_{S})\geq\ell_{\star}$,
$\frac{\rho}{4\sqrt{S}}\sum_{s=0}^{S-1}\mathbb{E}\|G_{k_{s}}\|_{F}^{2}\leq\mathbb{E}[\ell(\widetilde{W}_{0})]-\ell_{\star}+\frac{1}{\rho}\bar{E}_{\rm sq}+LC+A_{1}+A_{2}.$
Finally, since $\sum_{s=0}^{S-1}\mathbb{E}\|G_{k_{s}}\|_{F}^{2}\geq S\min_{s}\mathbb{E}\|G_{k_{s}}\|_{F}^{2}$, divide both sides by $\frac{\rho}{4}\sqrt{S}$ to get
$\min_{0\leq s\leq S-1}\mathbb{E}\|G_{k_{s}}\|_{F}^{2}\leq\frac{4}{\rho\sqrt{S}}\left(\mathbb{E}[\ell(\widetilde{W}_{0})]-\ell_{\star}+LC+A_{1}+A_{2}+\frac{1}{\rho}\bar{E}_{\rm sq}\right).$
This finishes the proof. ∎
Finally, we present the convergence rate in terms of the total number of iterations, as stated in Theorem 5.1, using the relationship that each restart cycle consists of $K$ iterations.
Corollary 1 (Convergence rate in total iterations).
Let $T$ be the total number of iterations allowed. Let $K$ be the fixed number of Adam steps per restart cycle, such that the number of restart cycles is $S=\lfloor T/K\rfloor$. Under the conditions of Theorem H.3, the convergence rate with respect to the minimum gradient norm over all $T$ iterates is:
$\min_{1\leq k\leq T}\mathbb{E}\|G_{k}\|_{F}^{2}\;\leq\;\frac{C_{0}}{\sqrt{\lfloor T/K\rfloor}}\;=\;\mathcal{O}\left(\frac{1}{\sqrt{T}}\right),$ (61)
where $C_{0}:=\frac{4}{\rho}\left(\mathbb{E}[\ell(\widetilde{W}_{0})]-\ell_{\star}+LC+(A_{1}+A_{2})+\frac{1}{\rho}\bar{E}_{\rm sq}\right)$.
Proof.
Let $\mathcal{S}=\{k_{0},k_{1},\dots,k_{S-1}\}$ be the set of indices corresponding to restart points. Since the set of restart points is a subset of all iterates $\{1,\dots,T\}$, the minimum over all iterates is upper bounded by the minimum over the restart points:
$\min_{1\leq k\leq T}\mathbb{E}\|G_{k}\|_{F}^{2}\;\leq\;\min_{s\in\{0,\dots,S-1\}}\mathbb{E}\|G_{k_{s}}\|_{F}^{2}.$
Substituting $S=\lfloor T/K\rfloor$ into the bound from Theorem H.3:
$\min_{0\leq s\leq S-1}\mathbb{E}\|G_{k_{s}}\|_{F}^{2}\leq\frac{C_{0}}{\sqrt{S}}=\frac{C_{0}}{\sqrt{\lfloor T/K\rfloor}}.$
For large $T$, we have $\sqrt{\lfloor T/K\rfloor}\approx\sqrt{T/K}$. Thus, the bound simplifies to:
$\frac{C_{0}}{\sqrt{T/K}}=\frac{C_{0}\sqrt{K}}{\sqrt{T}}.$
Thus, the rate for the iterates is $\mathcal{O}(1/\sqrt{T})$. ∎